### v2

# When Personas Override Payoffs: Role Identity Bias in Multi-Agent LLM Decision-Making

Abstract.
Large language models are increasingly deployed in multi-agent systems for strategic tasks, yet how design choices such as role-based personas and payoff visibility affect their reasoning capabilities remains poorly understood. We investigate whether multi-agent systems function as strategic reasoners capable of payoff optimization or as identity-driven actors that prioritize role alignment over explicit incentives. We use Nash equilibrium achievement in game-theoretic settings as a diagnostic tool to assess reasoning capabilities. We do so through systematic experiments with four Large language model (LLM) architectures (Qwen-7B/32B, Llama-8B, Mistral-7B) across complex environmental decision-making scenarios in four-agent strategic games. We demonstrate that role identity bias fundamentally alters strategic reasoning, i.e. when personas are active, all models achieve near-zero Nash equilibrium rates (0–6.7%) in scenarios where payoff-optimal equilibria exist, despite having complete payoff information. Removing personas and providing explicit payoffs enables Qwen models to achieve high Nash equilibrium rates (65–90%) in these scenarios, demonstrating that both conditions are necessary for strategic reasoning capabilities. Personas systematically bias equilibrium selection toward socially preferred outcomes, with 100% of Nash equilibria being Green Transition (socially preferred) when personas are present, while models completely fail to reach equilibrium in scenarios where Tragedy is optimal. The effect of explicit payoffs depends entirely on persona presence, revealing that design choices about personas and payoff visibility interact in ways that fundamentally determine system behavior. We observe striking model-dependent patterns where Qwen architectures show high sensitivity to both personas and payoff visibility, while Llama and Mistral maintain rigid reasoning patterns across conditions. These findings establish that representational choices are substantive governance decisions that determine whether multi-agent systems function as strategic reasoners or identity-driven actors, with important implications for deployment in strategic settings.
Large Language Models, Game Theory, Semantic Reasoning, Multi-Agent Systems, Environmental Decision-Making
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03-05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processingccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Model verification and validationccs: Computing methodologies Multi-agent systemsccs: Computing methodologies Cooperation and coordination
1. Introduction
Multi-agent systems powered by large language models are increasingly being deployed in policy simulation, economic modeling, negotiation platforms and other settings that require strategic reasoning (Merrill and Srivastava, 2025; Li et al., 2024; Bianchi et al., 2024a). In these applications, agents are frequently assigned role-based personas to represent different stakeholders or perspectives (Merrill and Srivastava, 2025; Xu et al., 2024b). These personas enable more realistic simulations and help systems model complex multi-stakeholder scenarios (Park et al., 2023), particularly in environmental and climate-related decision-making, where heterogeneous stakeholders interact strategically and coordination failures can lead to persistent social inefficiencies (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024). However, a critical question remains: when agents with role-based personas face scenarios that demand strategic decision making, do they follow explicit payoffs or their assigned identities?
This question determines whether multi-agent large language model (LLM) systems behave as strategic reasoners or as identity-driven actors who prioritize role alignment over payoff maximization (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025). If personas override explicit payoff information, a system designed for strategic analysis might instead follow the persona-consistent behavior that fails to optimize over encoded incentives (Neumann et al., 2025; Herr et al., 2024). This creates a fundamental tension between the representational goals of persona-based systems and their ability to reason strategically when payoff information is available (Bender et al., 2021; Duan et al., 2024b). Understanding how personas interact with payoff information in multi-agent settings is essential to make deployment decisions (Weidinger et al., 2022b; Chan et al., 2024; Casper et al., 2024; Rivera et al., 2024; Chan et al., 2024). However, this gap remains largely unexplored. Existing work has examined personas, payoff visibility, and model differences separately but has not systematically studied how these factors together determine system behavior in strategic settings (Neumann et al., 2025; Duan et al., 2024b; Akata et al., 2025; Huang et al., 2025; Sun et al., 2025; Zhang et al., 2024).
Refer to caption Flow diagram showing four agents with chain-of-thought reasoning making simultaneous action selections A flow diagram illustrating the multi-agent game structure with four agents (Environmentalist, Government, Industrialist, Citizen) engaging in simultaneous action selection. Each agent is represented by a robot figure with a dotted speech bubble above it containing chain-of-thought (CoT) reasoning text that shows how the agent analyzes the scenario and their role. The Environmentalist’s CoT begins with ”From the scenario, it is clear that…”, the Government’s with ”The scenario suggests that…”, the Industrialist’s with ”As an Industrialist, I need to…”, and the Citizen’s with ”Given my role as a citizen, I must…”. Each agent has two binary action choices (shown on arrows): Environmentalist chooses CAMPAIGN or NOCAMPAIGN, Government chooses REGULATE or NOREG, Industrialist chooses POLLUTE or CLEAN, and Citizen chooses SUPPORT_GREEN or BUY_CHEAP. All four arrows converge into a central ”Strategy Profile” box, representing the simultaneous action selection where all agents choose their actions at the same time, and the combination of all four actions forms the final strategy profile that is evaluated for Nash equilibrium.
Figure 1. The diagram shows our multi-agent game structure with chain-of-thought reasoning. Each agent analyzes the scenario and their role identity, then selects one of two binary actions. The arrows from each agent converge to a central Strategy Profile box, representing simultaneous action selection where all four agents choose actions concurrently, and their combined choices form the strategy profile evaluated for Nash equilibrium. The dotted speech bubbles contain each agent’s chain-of-thought reasoning process.
Understanding how role identity and payoff optimization interact is critical in multi-agent strategic games. We study environmental policy scenarios with four agents (Industrialist, Government, Environmental Activist, and Citizen) coordinating to reach Nash equilibria (Nash, 1951), which can yield either Green Transition (socially preferred and environmentally sustainable) or Tragedy of Commons (payoff-optimal but environmentally damaging) outcomes (Hardin, 1968). Through a $\times2$ factorial experiment manipulating persona presence and payoff visibility across four model architectures, we examine whether agents prioritize role-aligned actions or payoff-optimal strategies. Our findings show that personas fundamentally alter strategic reasoning, preventing payoff optimization: when personas are active, all models achieve near-zero Nash equilibrium rates in economic scenarios where Tragedy of Commons is payoff-optimal, despite complete payoff information. Removing personas and providing explicit payoffs enables certain models to achieve high Nash equilibrium in these scenarios, demonstrating both conditions are necessary for strategic reasoning. Personas systematically bias equilibrium selection toward socially preferred outcomes, with Nash equilibria being exclusively Green Transition when personas are present, while models fail to reach equilibrium in scenarios where Tragedy is optimal. The effect of explicit payoffs depends entirely on persona presence, and we observe model-dependent patterns where some architectures show high sensitivity to both personas and payoff visibility, while others maintain rigid reasoning patterns across conditions. This work provides key contributions to understanding LLM behavior in multi-agent strategic settings. We demonstrate that role identity bias completely overrides explicit payoff optimization, preventing Nash equilibrium. Additionally, personas and payoff visibility interact in ways that fundamentally determine system behavior, meaning design choices about representation formats are substantive governance decisions. We also reveal model-dependent persona effects, suggesting different architectures implement strategic reasoning through different mechanisms. These findings have practical implications for deployment, as personas alter system behavior in ways that may not be immediately apparent to users, showing that deploying personas requires careful consideration of how identity interacts with payoff optimization in strategic decision making.
2. Related Work
2.1. LLMs in Multi-Agent Systems
Large language models have increasingly been studied as agents in multi-agent systems, with applications spanning negotiation, deliberation, debate, and collaborative problem-solving (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Tran et al., 2025; Guo et al., 2024; Xu et al., 2024a; Agashe et al., 2025; Bonabeau, 2002; Epstein and Axtell, 1996). Early work demonstrated that LLMs can sustain multi-turn interactions and coordinate actions through natural language communication while exhibiting emergent collective behaviors (Xu et al., 2024a; Bo et al., 2024). Following studies explored deliberative and collaborative agent frameworks, emphasizing consensus formation and group-level reasoning (Abdelnabi et al., 2024; Bianchi et al., 2024b; Agashe et al., 2025; Bo et al., 2024).
Beyond task completion, multi-agent LLM systems are increasingly used as agent-based simulations to study complex social, economic, and institutional dynamics, reflecting a broader resurgence of agent-based modeling for analyzing emergent behavior in policy and social systems (Bonabeau, 2002; Epstein and Axtell, 1996; Park et al., 2022; Aher et al., 2023; Zheng et al., 2020; Gao et al., 2024).
However, most of this literature focuses on task-oriented collaboration or collective sense-making, rather than strategic settings in which agents must reason about incentives, anticipate others’ actions, and converge to equilibria (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b). As a result, core game-theoretic questions such as payoff optimization, equilibrium selection, and incentive compatibility remain underexplored in multi-agent LLM research.
2.2. LLMs and Game-Theoretic Reasoning
Classical game theory models agents as rational payoff maximizers who compute Nash equilibria under complete information (Nash, 1951). These game-theoretic models have been particularly influential in environmental decision making, where multi-actor coordination failures give rise to socially suboptimal outcomes such as over-extraction and pollution, while emphasizing the biospheric and socio-technical limits of purely objective-driven optimization (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024; Domínguez Hernández et al., 2024). A growing body of work has examined whether LLMs approximate such behavior in strategic games, often focusing on simple two-player settings such as Prisoner’s Dilemma, coordination games, or public goods games (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024; Fish et al., 2025). These studies consistently show that LLMs exhibit cooperative or norm-aligned behavior even when defection or self-interested actions are payoff-dominant (Akata et al., 2025; Fan et al., 2023; Herr et al., 2024). Several works suggest that apparent equilibrium behavior in LLMs may arise through semantic or narrative heuristics rather than explicit optimization over incentives (Herr et al., 2024; Macmillan-Scott and Musolesi, 2024; Fish et al., 2025; Bender et al., 2021). When payoffs are implicit, models can sometimes reach Nash equilibria “by accident,” but this behavior is unstable and sensitive to framing. Existing studies rarely isolate the effect of payoff visibility or systematically compare implicit and explicit payoff representations, particularly in multi-agent settings beyond two players (Huang et al., 2025; Duan et al., 2024b; Sun et al., 2025; Duan et al., 2024a). As a result, it remains unclear whether explicit payoff matrices enable principled game-theoretic reasoning or whether LLMs continue to rely on semantic shortcuts even when incentives are fully specified.
2.3. Role Personas and Identity Bias
Role-based prompting and persona conditioning are widely used to steer LLM behavior, with prior work demonstrating effects in persuasion, negotiation, and debate (Abdelnabi et al., 2024; Bianchi et al., 2024b; Hua et al., 2024; Wang et al., 2025; Liang et al., 2024; Neumann et al., 2025). Personas are often introduced to improve realism or alignment with human expectations and are commonly treated as stylistic or surface-level interventions (Park et al., 2023, 2022). However, more recent work has identified a stronger phenomenon commonly referred to as role identity bias, which refers to the tendency of LLMs to adhere to persona-consistent actions even when they conflict with explicit goals, instructions, or incentives, reflecting broader fairness and accountability concerns about how institutional roles and power are encoded in algorithmic systems (Neumann et al., 2025; Macmillan-Scott and Musolesi, 2024; Aher et al., 2023; Rakova and Dobbe, 2023). For example, agents assigned roles associated with profit maximization or authority may favor actions consistent with those identities despite countervailing information (Neumann et al., 2025). Prior studies largely focus on single-agent or two-player interactions, leaving open how role identity bias manifests in larger strategic systems and whether it interferes with equilibrium convergence when explicit payoff information is available (Akata et al., 2025; Herr et al., 2024; Huang et al., 2025). These questions are especially consequential in environmental policy settings, where agents are explicitly cast as stakeholders with conflicting priorities, and where equilibrium selection determines whether coordination or collective-action failure emerges (Ostrom, 1990; Barrett, 2003).
2.4. Information Presentation and Reasoning Fidelity
A considerable amount of literature examines how information presentation affects LLM reasoning across domains, including mathematical problem solving, logical inference, and strategic decision-making, emphasizing that linguistic framing and role description can materially alter behavior even when underlying information is unchanged (Wei et al., 2022; Duan et al., 2024a, b; Herr et al., 2024; Kim et al., 2024; Neumann et al., 2025; Bender et al., 2021; Blodgett et al., 2020). Structured representations, explicit constraints, and intermediate reasoning steps often improve performance, but gains are inconsistent and highly task-dependent (Wei et al., 2022; Duan et al., 2024a, b). Chain-of-thought (CoT) prompting has been shown to improve reasoning accuracy by encouraging intermediate decomposition (Wei et al., 2022). Most CoT research focuses on single-agent tasks, with limited investigation into multi-agent strategic interaction (Duan et al., 2024b, a; Huang et al., 2025). Moreover, recent work cautions that improved surface-level reasoning does not necessarily imply faithful optimization or robust decision-making (Casper et al., 2024; Kim et al., 2024; Weidinger et al., 2022a). The relative importance of payoff visibility versus reasoning scaffolds in achieving equilibrium behavior therefore remains insufficiently understood (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025).
2.5. Model and Architectural Effects in Strategic Behavior
Differences in model architecture and training methods have been shown to affect LLM behavior in strategic and social reasoning tasks (Akata et al., 2025; Herr et al., 2024; Huang et al., 2025; Duan et al., 2024b). Model scale alone does not reliably predict performance in strategic settings, and different model families exhibit distinct inductive biases and behavioral tendencies (Akata et al., 2025; Herr et al., 2024; Duan et al., 2024b). Comparative studies suggest that some models rely more heavily on semantic cues and role-based associations, while others appear less sensitive to persona framing (Neumann et al., 2025; Macmillan-Scott and Musolesi, 2024). These findings imply that strategic reasoning in LLMs may be implemented through qualitatively different mechanisms across architectures, with important implications for robustness and deployment in high-stakes settings (Weidinger et al., 2022a; Casper et al., 2024).
Taken together, these strands of work point to an incomplete understanding of strategic behavior in multi-agent LLM systems. Prior work has examined collaboration, game-theoretic reasoning, personas, and information presentation largely in isolation (Guo et al., 2024; Tran et al., 2025; Sun et al., 2025; Akata et al., 2025; Neumann et al., 2025; Kim et al., 2024), but prior existing work has not systematically studied how these factors interact in multi-agent settings with explicit payoff optimization (Sun et al., 2025; Huang et al., 2025; Duan et al., 2024b; Fish et al., 2025). Specifically, it remains unclear whether explicit payoff information enables true game-theoretic reasoning, how role identity bias interferes with equilibrium convergence, and whether these effects vary across model architectures (Fan et al., 2023; Herr et al., 2024; Neumann et al., 2025; Akata et al., 2025). Our work addresses this gap by jointly analyzing role identity, payoff visibility, and reasoning scaffolds in four-agent strategic games, revealing systematic failures of payoff-optimal reasoning and model-dependent patterns of equilibrium selection (Sun et al., 2025; Duan et al., 2024b; Casper et al., 2024).
3. Experimental Design
Our experiment design focuses on distinguishing between identity-driven behavior and payoff-optimal reasoning. This required two key decisions: selecting an appropriate domain that would create a testable tension between role identity bias and strategic payoff maximization, and developing an experimental design that could systematically test how these mechanisms interact in LLM decision-making. We first introduce the domain we chose and explain why it is well-suited for this research, then describe the game structure and experimental design.
3.1. Domain Choice
We chose environmental decision-making as our experimental domain because it creates a natural tension between semantic associations and strategic optimization. Environmental policy involves multiple stakeholders (industry, government, activists, and citizens) with clearly defined but often conflicting interests that map directly to game-theoretic player types. This enables a 4-player game structure that captures real-world dynamics.
Most critically, role-based personas create a conflict between identity-driven behavior and payoff-optimal reasoning. Actions aligned with positive semantic associations from training data may not be payoff-optimal, particularly in economic scenarios where Nash equilibrium may conflict with these associations. In economic scenarios, the Nash equilibrium that maximizes payoffs may involve actions that conflict with these positive associations. This creates a natural experimental setup. If role identity bias dominates, models choose persona-aligned actions, and if game-theoretic reasoning dominates, models choose payoff-optimal actions even when they conflict with role identity. By systematically manipulating persona presence and payoff visibility, we determine whether role identity bias prevents payoff-optimal equilibria and under what conditions payoff-optimal reasoning can be recovered. Our experiments use a 4-player environmental decision-making strategy game where LLM agents play different roles, as shown in Fig. 1:
  1. (1)
    Industrialist (I): chooses between POLLUTE, representing profit-maximizing production with minimal environmental safeguards, and CLEAN, representing investment in environmentally friendly production at higher short-term cost.
  2. (2)
    Government (G): chooses between NOREG, maintaining the regulatory status quo, and REGULATE, imposing environmental regulations that incur enforcement and political costs.
  3. (3)
    Environmental Activist (A): chooses between NOCAMPAIGN, conserving resources and remaining inactive, and CAMPAIGN, mobilizing public pressure through advocacy and protest.
  4. (4)
    Citizen/Consumer (C): chooses between BUY_CHEAP, prioritizing lower-cost goods and short-term affordability, and SUPPORT_GREEN, accepting higher costs to support environmentally sustainable options.
These four roles were chosen to capture the essential stakeholders in environmental policy decisions. The Industrialist represents private industry facing the trade-off between profit and environmental responsibility. The Government represents regulatory authority balancing economic growth with environmental protection. The Environmental Activist represents advocacy groups pushing for environmental action regardless of economic costs. The Citizen represents the general public facing the tension between environmental values and economic constraints. This structure allows us to distinguish between economic actors (Industrialist and Government) and social actors (Activist and Citizen).
3.1.1. Strategy Profiles and Nash Equilibrium
A strategy profile is a 4-tuple specifying each player’s action, for example ‘(Clean, Regulate, Campaign, SupportGreen)‘. Two profiles are particularly important:
  • Green Transition: ‘(Clean, Regulate, Campaign, SupportGreen)‘ - All players choose environmentally-friendly actions
  • Tragedy of Commons: ‘(Pollute, NoReg, NoCampaign, BuyCheap)‘ - All players prioritize self-interest over environmental concerns
After all agents make their decisions simultaneously, we compute whether the resulting strategy profile is a Nash equilibrium. A strategy profile is a Nash equilibrium if no player can improve their payoff by unilaterally changing their action, given the payoff structure and other players’ actions. If models are using game-theoretic reasoning, they should converge to Nash equilibria. If role identity bias dominates, they may choose profiles that align with their respective persona’s expectations but are not Nash equilibria.
The payoff structure determines which profiles are Nash equilibria. For example, if payoffs are structured so that Green Transition gives the highest payoffs to all players, then Green Transition is a Nash equilibrium. If payoffs are modified so that polluting is more profitable, then Tragedy of Commons might become the Nash equilibrium instead.
3.1.2. Scenarios
We tested 53 scenarios across three scenario types, all designed with payoff-coupled narratives where the scenario description aligns with the underlying payoff structure. All scenario narratives were generated using GPT-5-mini [^1] Azure OpenAI GPT-5-mini (model version 2025-08-07) and subsequently reviewed and filtered to ensure consistency with the intended payoff structure and equilibrium class [^2] Model-generated scenarios were used to reduce author-specific framing bias and to enable systematic variation across narrative surface forms while holding the underlying game structure fixed. Scenarios were retained only if the narrative description unambiguously aligned with the corresponding payoff incentives, and no model was evaluated on scenarios it had previously generated.. Models are exposed to each scenario in two versions: one where payoffs are implicit and must be inferred from the narrative description, and one where explicit payoff matrices are provided. Scenarios are categorized into three subtypes  [^3] Example from the scenario set and a complete list of scenario subtypes are included in Appendix A.7.:
  • Economic/Tragedy Scenarios: These scenarios describe economic crises where the payoff structure makes Tragedy of Commons the dominant Nash equilibrium.
  • Environmental/Green Scenarios: These scenarios describe environmental emergencies where the payoff structure makes Green Transition the dominant Nash equilibrium.
  • Mixed/Complex Scenarios: These scenarios involve situations where economic and environmental considerations are intertwined, creating complex payoff structures that may not clearly favor either Green Transition or Tragedy of Commons as the dominant equilibrium.
The scenario set is intentionally imbalanced across these two classes. Environmental scenarios are overrepresented because they reflect a primary real-world domain in which role-based multi-agent LLM systems are actively deployed for policy analysis, sustainability planning, and climate governance. Economic Tragedy-of-the-Commons scenarios serve as contrast cases, enabling controlled comparison between settings where payoff-optimal actions conflict with positive semantic associations and settings where payoff-optimal actions align with them. This design allows us to test how models balance identity-driven behavior and payoff-optimal reasoning, whether they can infer payoffs from narrative descriptions, whether explicit payoff information enables payoff-optimal reasoning, and how role identity bias interacts with payoff visibility to affect equilibrium selection.
Table 1. Overall Nash Equilibrium Rates: With Personas vs Without Personas. This table compares overall Nash equilibrium rates across all scenarios (economic + environmental) between conditions with and without personas. The “Difference” column shows the change in percentage points when personas are removed (negative values indicate personas help, positive values indicate personas hurt).
Model Payoffs With Persona Without Persona Difference
Llama-8B Hidden 77.4% (205/265) 75.5% (200/265) -1.9 pp
Llama-8B Visible 77.0% (204/265) 72.1% (191/265) -4.9 pp
Mistral-7B Hidden 77.4% (205/265) 75.8% (201/265) -1.6 pp
Mistral-7B Visible 69.8% (185/265) 38.9% (103/265) -30.9 pp
Qwen-32B Hidden 69.1% (183/265) 74.3% (197/265) +5.2 pp
Qwen-32B Visible 57.7% (153/265) 33.2% (88/265) -24.5 pp
Qwen-7B Hidden 75.5% (200/265) 75.5% (200/265) 0.0 pp
Qwen-7B Visible 75.5% (200/265) 55.5% (147/265) -20.0 pp
3.2. Chain-of-Thought Reasoning Analysis
We analyze the chain-of-thought reasoning generated by models during action selection to understand how role identity bias manifests in decision-making processes. For each agent’s action selection, we extract the rationale field containing the model’s reasoning and analyze keyword patterns that distinguish different reasoning mechanisms (see Appendix C for keyword categories). We compare keyword frequency distributions across experimental conditions to identify how personas and payoff visibility affect reasoning style. From this analysis, we infer whether models are using game-theoretic reasoning (direct payoff optimization) or identity-driven reasoning (role-aligned decision-making). This allows us to understand the mechanisms behind observed Nash equilibrium patterns and how role identity bias shifts reasoning from payoff-optimal to identity-aligned strategies. While we reference key CoT findings in the main text to support our quantitative results, we do not analyze these results in detail here. A comprehensive analysis of keyword frequency patterns, reasoning shifts, and detailed visualizations are presented in the Appendix (Figures 5-8).
3.3. Models and Experimental Conditions
Our experimental design follows a $\times2$ factorial structure, crossing role identity and payoff visibility, and tests 53 environmental decision-making scenarios in a 4-agent simultaneous-move game. We use four instruction-tuned, open-source LLMs: Qwen2.5-7B (Qwen et al., 2025), Qwen2.5-32B (Qwen et al., 2025), Llama-3.1-8B (Grattafiori et al., 2024), and Mistral-7B (Jiang et al., 2023). For each model, condition, and scenario, a single model instance takes on all four agent roles (Industrialist, Government, Environmental Activist, Citizen) and is prompted independently for each role. Agents select actions in a fixed order [^4] This order is fixed (not randomized) to ensure reproducibility across experiments, and each configuration is repeated five times.
We compare two persona conditions: with personas (role-specific descriptions reflecting stakeholder motivations) and without personas (agents instructed only on Nash equilibrium reasoning). We also compare two payoff conditions: hidden (agents infer incentives from scenario descriptions) and visible (agents receive explicit payoff matrices). All experiments use chain-of-thought reasoning instructions. We evaluate outcomes along two dimensions: whether the resulting action profile constitutes a Nash equilibrium and which canonical equilibrium is selected (Green Transition or Tragedy of Commons).
4. Results
We now present quantitative results organized by the research questions that we aim to answer. We report results by answering the five research questions ,across the settings mentioned above, based on our experiments backed by tables and figures.
RQ.1. Does Role Identity Bias Prevent Nash Equilibrium Achievement in Economic Scenarios?
Answer: Yes. When agents are given personas, 12 out of 16 experiments across different models and variants achieve 0% Nash equilibrium in economic scenarios despite explicit Nash equilibrium instructions and chain-of-thought reasoning. Fig. 3 shows our findings. When payoff matrices are visible with personas, all models achieve 0% Nash in economic scenarios, confirming that role identity bias overrides explicit payoff information. Models select Green Transition actions (20-65%, Fig. 2 shows detailed Green Transition selection rates) in economic scenarios when personas are present, even when Tragedy is payoff-dominant, demonstrating role identity bias overriding payoff optimization.
Table 2. Environmental Scenario Nash Equilibrium Rates: With Personas vs. Without Personas. This table shows Nash equilibrium rates in environmental scenarios (where Green Transition is the payoff-optimal equilibrium) between conditions with and without personas.
Model Payoffs With Persona Without Persona Difference
Llama-8B Hidden 100.0% (205/205) 97.6% (200/205) -2.4 pp
Llama-8B Visible 99.5% (204/205) 93.2% (191/205) -6.3 pp
Mistral-7B Hidden 100.0% (205/205) 98.0% (201/205) -2.0 pp
Qwen 32B Hidden CoT 89.3% (183/205) 96.1% (197/205) +6.8 pp
Qwen 7B Hidden CoT 95.6% (196/205) 95.6% (196/205) 0.0 pp
Table 3. Economic Scenario Nash Equilibrium Rates: Hidden vs Visible Payoffs. This table compares Nash equilibrium rates in economic scenarios (where Tragedy is payoff-optimal) between Hidden and Visible payoff conditions. Each row shows a different model-persona combination.
Model Persona Hidden Visible Difference
Llama-8B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
Llama-8B No 0.0% (0/60) 0.0% (0/60) 0.0 pp
Mistral-7B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
Mistral-7B No 0.0% (0/60) 0.0% (0/60) 0.0 pp
Qwen-32B Yes 0.0% (0/60) 0.0% (0/60) 0.0 pp
Qwen-32B No 0.0% (0/60) 90.0% (54/60) +90.0 pp
Qwen-7B Yes 6.7% (4/60) 0.0% (0/60) -6.7 pp
Qwen-7B No 6.7% (4/60) 65.0% (39/60) +58.3 pp
Our findings show that role identity bias prevents Nash equilibrium when personas are present, causing models to align actions with persona expectations rather than payoff-optimal Tragedy actions. Chain-of-thought analysis reveals that with personas, models show more “short-term” and “social-moral” reasoning patterns, indicating identity-driven rather than payoff-optimal decision-making. CoT analysis also reveals different reasoning patterns between models. Llama and Mistral show consistent patterns regardless of payoff visibility, while Qwen models show some payoff-focused keywords but still cannot overcome identity bias to select Tragedy equilibrium.
RQ.2. How does removing personas affect Nash equilibrium achievement?
Answer: Removing personas enables partial Tragedy selection in economic scenarios but reduces overall Nash rates by 1.9-7.3 percentage points. Table 1 compares overall Nash equilibrium rates across all scenarios (economic and environmental) between conditions with and without personas. This table demonstrates that removing personas generally reduces Nash rates for most models, except Qwen-32B with hidden payoffs, which shows a slight increase.
Personas both enable and constrain Nash equilibrium achievement. Overall Nash rates decrease by 1.9-4.9 pp for Llama and by 1.6 pp for Mistral when personas are removed, and environmental scenario Nash rates decrease by 2-6 pp (Table 2) without personas for most models, indicating personas enable better coordination in Green Transition scenarios. CoT analysis shows that with personas, models exhibit more “social-moral” reasoning, which aligns with socially preferred Green Transition outcomes. However, Qwen-32B shows the opposite pattern, where Nash rates increase by 5.2 pp (Table 1) without personas, suggesting personas constrain Qwen’s ability to reach Nash equilibrium. CoT analysis reveals that removing personas shifts reasoning from “social-moral” and “short-term” patterns to “strategic” and “long-term” thinking for Qwen models, indicating a shift from identity-driven to payoff-optimal reasoning.
In economic scenarios, removing personas reduces Green Transition selection by 10-37 percentage points (see Table 10 in the Appendix), indicating partial Tragedy selection when personas are absent. However, removing personas alone is insufficient. Without personas but with hidden payoffs, all models still achieve 0% Nash in economic scenarios. Only Qwen models with visible payoffs and no personas can select Tragedy equilibrium (65-90% Nash, Fig. 3), demonstrating that both removing personas and providing explicit payoffs are necessary for payoff-optimal reasoning.
RQ.3. Does information presentation format (hidden vs visible payoffs) affect Nash equilibrium achievement?
Answer: Yes. Hidden payoffs achieve similar or higher Nash rates than visible payoffs when personas are present, but visible payoffs enable Tragedy equilibrium selection when personas are removed (for Qwen models).
Refer to caption Grouped bar chart showing Green Transition action selection rates across four models and four experimental conditions A grouped bar chart displaying Green Transition action selection rates (percentage) in economic scenarios across four language models (Qwen-7B, Qwen-32B, Llama-8B, Mistral-7B) and four experimental conditions: Personas with Visible Payoffs, Personas with Hidden Payoffs, No Personas with Visible Payoffs, and No Personas with Hidden Payoffs. The x-axis shows the four condition combinations, and the y-axis shows the Green Transition selection rate from 0% to 100%. Each model is represented by a distinct color and pattern: Qwen-7B uses solid blue bars, Qwen-32B uses light blue bars with dotted pattern, Llama-8B uses orange bars with diagonal lines, and Mistral-7B uses light orange bars with crosshatch pattern. The bars are grouped by condition, allowing comparison of each model’s Green Transition selection rate across different experimental settings. This visualization demonstrates how often models select Green Transition actions (socially preferred but payoff-suboptimal in economic scenarios) rather than Tragedy of Commons actions (payoff-optimal), revealing the extent of role identity bias across different model architectures and experimental conditions.
Figure 2. Green Transition Action Selection Rates in Economic Scenarios. This grouped bar chart shows the percentage of action profiles where models select Green Transition actions (socially preferred but payoff-suboptimal) in economic scenarios where the Tragedy of Commons is the payoff-optimal equilibrium.
Refer to caption Heatmap showing Nash equilibrium rates across four models and four condition combinations A heatmap displaying Nash equilibrium rates (0-100%) for economic scenarios across four language models (Llama 8B, Mistral 7B, Qwen 32B, Qwen 7B) and four experimental conditions: Hidden Payoff with Persona, Hidden Payoff without Persona, Visible Payoff with Persona, and Visible Payoff without Persona. The heatmap uses a blue color gradient where lighter shades represent 0% Nash rates and darker shades represent higher rates. Most cells show 0.0% (light blue), with two notable exceptions: Qwen 32B in the Visible Payoff without Persona condition shows 90.0% (dark blue), and Qwen 7B in the same condition shows 65.0% (medium-dark blue). This visualization clearly demonstrates that only Qwen models achieve high Nash equilibrium rates in economic scenarios, and only when both conditions are met: payoffs are visible and personas are removed. All other model-condition combinations result in 0% Nash rates, with Qwen 7B showing minor exceptions (6.7%) in Hidden Payoff conditions regardless of persona presence.
Figure 3. Economic Scenarios Nash Equilibrium Rates by Model and Condition. This heatmap visualizes Nash equilibrium achievement rates in economic scenarios where Tragedy of Commons is the payoff-optimal equilibrium across all four models and four experimental conditions. The color intensity represents the Nash rate percentage, with darker blue indicating higher rates.
Table 4. Equilibrium Selection Patterns: With Personas Condition. Note: “Green % of Nash” shows the percentage of Nash equilibria that are Green Transition. The denominators vary because different models/variants achieve Nash equilibrium at different rates (see “Overall Nash” column), but when Nash is achieved, it is always Green Transition (100%).
Model Variant Overall Nash Green Transition Rate Tragedy Rate Green % of Nash
Llama-8B Hidden 77.4% (205/265) 87.5% (232/265) 0.0% (0/265) 100.0% (205/205)
Llama-8B Visible 77.0% (204/265) 77.0% (204/265) 0.0% (0/265) 100.0% (204/204)
Mistral-7B Hidden 77.4% (205/265) 90.6% (240/265) 0.0% (0/265) 100.0% (205/205)
Mistral-7B Visible 69.8% (185/265) 69.8% (185/265) 0.0% (0/265) 100.0% (185/185)
Qwen-32B Hidden 69.1% (183/265) 69.8% (185/265) 0.0% (0/265) 100.0% (183/183)
Qwen-32B Visible 57.7% (153/265) 57.7% (153/265) 0.0% (0/265) 100.0% (153/153)
Qwen-7B Hidden 75.5% (200/265) 75.5% (200/265) 0.0% (0/265) 100.0% (200/200)
Qwen-7B Visible 75.5% (200/265) 75.5% (200/265) 0.0% (0/265) 100.0% (200/200)
Table 11 in the Appendix compares overall Nash equilibrium rates between hidden and visible payoff conditions. When personas are present, hidden payoffs achieve equal or higher Nash rates than visible payoffs for all models. This result suggests that when role identity bias is present, inferring payoffs from scenario descriptions supports better coordination than explicit payoff matrices. CoT analysis confirms this claim. With personas, hidden payoffs show more “inferential” reasoning (inferring payoffs from scenarios), while visible payoffs show more “explicit_info” keywords, but both achieve similar Nash rates because persona expectations align with scenario descriptions.
This pattern reverses when personas are removed. Table 3 shows that visible payoffs enable much higher Nash rates than hidden payoffs for Qwen models when personas are absent (+58.3 to +90.0 pp difference). CoT analysis reveals that without personas, visible payoffs shift Qwen models to nearly universal “game-theoretic” and “payoff-focused” reasoning (99.7% of rationales for Qwen-32B), with “strategic” keywords appearing 368 times, demonstrating direct payoff optimization. This demonstrates that explicit payoff information enables game-theoretic reasoning when role identity bias is removed. However, this capability is model-dependent. Only Qwen models show this behavior, while Llama and Mistral achieve 0% Nash even without personas, suggesting architectural differences in how models process explicit payoff information.
The effect of visible payoffs also differs across scenario types. Table 5 shows that visible payoffs with no personas cause large decreases in environmental Nash for Qwen models, dropping from 96% to 17% for Qwen 32B, suggesting a trade-off where models select Tragedy in economic scenarios but fail to coordinate on Green in environmental scenarios. Mistral shows the opposite pattern, with visible payoffs and no personas, reducing environmental Nash from 98% to 50% (Table 5), while economic Nash remains at 0% (Table 3), suggesting Mistral cannot leverage explicit payoffs effectively regardless of scenario type.
Table 5. Environmental Scenario Nash Equilibrium Rates: Comparison of Hidden vs. Visible payoffs. Negative values indicate visible payoffs hurt, positive values indicate visible payoffs help.
Model Persona Hidden Visible Difference
Llama-8B Yes 100.0% (205/205) 99.5% (204/205) -0.5 pp
Llama-8B No 97.6% (200/205) 93.2% (191/205) -4.4 pp
Mistral-7B Yes 100.0% (205/205) 90.2% (185/205) -9.8 pp
Mistral-7B No 98.0% (201/205) 50.2% (103/205) -47.8 pp
Qwen 32B Yes 89.3% (183/205) 74.6% (153/205) -14.7 pp
Qwen 32B No 96.1% (197/205) 16.6% (34/205) -79.5 pp
Qwen 7B Yes 95.6% (196/205) 97.6% (200/205) +2.0 pp
Qwen 7B No 95.6% (196/205) 52.7% (108/205) -42.9 pp
RQ.4. Are persona effects model-dependent?
Answer: Yes. Qwen models show stronger persona dependence (9.7-10.0 pp average reduction in Nash rates without personas), while Llama-8B shows minimal dependence (3.4 pp average, range 1.9-4.9 pp). [^5] See Tables 6(b)6(a), and 12 in the Appendix for detailed persona effect statistics..
Table 6(a) quantifies the magnitude of persona effects on overall Nash equilibrium rates. All models show negative persona effects in the visible payoffs variant (-5 to -31 pp), but the mechanisms differ.
For Qwen models, removing personas enables payoff-optimal Tragedy equilibrium in economic scenarios but reduces Green Transition equilibrium in environmental scenarios when visible payoffs are present (96.1% $\to$ 16.6% Nash for Qwen-32B, Table 5), resulting in net negative overall effects (-20-25 pp, Table 6(a)). For Llama/Mistral, removing personas only reduces environmental Nash equilibrium without enabling economic Nash (0% remains), reflecting pure Nash rate loss. Llama-8B shows minimal persona dependence (1.9-4.9 pp reduction), while Mistral-7B shows large effects (-30.9 pp) due to reduced environmental Nash rates.
Analyzing the Chain-Of-Thought of the models shows how differently they reason. Llama maintains consistent reasoning patterns across conditions (minimal keyword differences), suggesting its reasoning is relatively unaffected by persona presence. Mistral shows payoff optimization keywords but cannot identify Tragedy equilibrium regardless of condition. Qwen models show dramatic shifts with and without personas. With personas, they exhibit “short-term” reasoning; without personas and with visible payoffs, they shift to “strategic” (368 mentions for Qwen-32B) and “long-term” (283 mentions) reasoning. 99.7% of the rationales [^6] See Section E for detailed COT analysis contain game-theoretic keywords, demonstrating a complete shift from identity-driven to payoff-optimal reasoning.
Additionally, model size does not predict persona dependence. Both Qwen-7B and Qwen-32B show similar reasoning pattern shifts, but Qwen-32B achieves higher Nash rates due to stronger payoff optimization.
Refer to caption Two-panel stacked bar chart showing Green Transition and Tragedy of Commons percentages A two-panel stacked bar chart displaying the percentage of Nash equilibria that are Green Transition versus Tragedy of Commons across four language models (Llama 8B, Mistral 7B, Qwen 32B, Qwen 7B) in the without personas condition. The left panel shows the Hidden Payoff condition, where all models achieve 100% Green Transition (blue bars with diagonal stripes). The right panel shows the Visible Payoff condition, where Llama 8B and Mistral 7B maintain 100% Green Transition, while Qwen 32B shows 11.4% Green Transition (green) and 88.6% Tragedy of Commons (orange-red), and Qwen 7B shows 89.8% Green Transition and 10.2% Tragedy of Commons. Each bar segment is labeled with its percentage value in white text. The legend is positioned below the figure, showing Green Transition in blue with diagonal stripes and Tragedy of Commons in orange-red with dots.
Figure 4. Equilibrium Selection Patterns: Green Transition vs Tragedy of Commons (Without Personas Condition). This figure shows the distribution of Nash equilibria between Green Transition and Tragedy of Commons outcomes across all four models in the no-personas condition.
RQ.5. Do personas enable equilibrium selection beyond pure payoff maximization?
Answer: Yes. When personas are present, 100% of Nash equilibria are Green Transition (socially preferred), showing that personas shift equilibrium selection from payoff-based to socially biased. Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants. Fig. 4 reveals model-dependent patterns when personas are removed. For Qwen-32B with visible payoffs, only 11.4% of Nash equilibria are Green Transition (10/88), meaning 88.6% are Tragedy, demonstrating complete reversal from the persona condition where 100% were Green Transition. For Qwen-7B with visible payoffs, 89.8% of Nash equilibria are Green Transition (132/147), but 65% Nash in economic scenarios 3 means most economic Nash is Tragedy. For Llama and Mistral, 100% of Nash equilibria remain Green Transition even without personas, indicating they cannot identify Tragedy equilibrium regardless of condition.
Table 6. Persona Effects on Nash Equilibrium Rates: Overall (left) and Economic Scenarios (right). Left table shows overall persona effects across all scenarios (economic and environmental combined). Right table shows persona effects specifically for economic scenarios (where Tragedy is payoff-optimal). Negative values (percentage points) indicate personas help and positive values (percentage points) indicate personas hurt.
(a) Overall Nash Equilibrium Rates and their dependence on personas for various LLM architectures.
Model Hidden CoT Visible CoT Average
Llama 8B -1.9 pp -4.9 pp -3.4 pp
Mistral 7B -1.6 pp -30.9 pp -16.3 pp
Qwen 32B +5.2 pp -24.5 pp -9.7 pp
Qwen 7B 0.0 pp -20.0 pp -10.0 pp
(b) Economic Scenarios
Model Hidden CoT Visible CoT Average
Llama 8B 0.0 pp 0.0 pp 0.0 pp
Mistral 7B 0.0 pp 0.0 pp 0.0 pp
Qwen 32B 0.0 pp +90.0 pp +45.0 pp
Qwen 7B 0.0 pp +65.0 pp +32.5 pp
CoT analysis confirms the mechanism. With personas, models show “social-moral” and “short-term” reasoning, indicating identity-driven decision-making. Without personas and with visible payoffs, Qwen models shift to nearly universal “game-theoretic” reasoning (99.7% of rationales for Qwen-32B), with “strategic” and “long-term” keywords unique to this condition, demonstrating a complete shift from identity-driven to payoff-optimal reasoning that enables equilibrium selection reversal. Llama and Mistral maintain consistent reasoning patterns regardless of condition, explaining their inability to shift equilibrium selection.
5. Discussion
We ground our discussion in specific empirical findings from above and discuss what these results imply for developers, practitioners, and system design.
5.0.1. Role Identity Bias Fundamentally Alters Strategic Reasoning
Personas prevent models from identifying payoff-optimal equilibria in economic scenarios, with 12 out of 16 experiments achieving 0% Nash equilibrium despite explicit payoff information (Fig. 3). Persona expectations override explicit incentive structures, leading to reasoning patterns shifting from strategic optimization to role-aligned decision-making. This demonstrates that personas function as strong normative constraints that can completely suppress game-theoretic reasoning capabilities, consistent with prior findings that role prompts and system-level personas systematically bias strategic decision-making in LLMs (Neumann et al., 2025; Herr et al., 2024).
Our results show that personas are not just cosmetic enhancements but design choices that fundamentally alter system behavior. Personas systematically bias equilibrium selection toward socially preferred outcomes (100% Green Transition with personas, Table 4). Developers should explicitly document and disclose persona choices, as they determine system behavior. Users require clear information about system behavior, particularly in applications such as policy analysis or economic simulation where they may assume agents are behaving optimally according to their payoffs.
5.0.2. Payoff-Optimal Reasoning Requires Removing Personas and Providing Explicit Payoffs
Our experiments show that strategic reasoning emerges only when both role identity bias is removed and explicit payoff information is provided simultaneously (Fig. 3). All models achieve 0% Nash in economic scenarios without personas but with hidden payoffs. Only Qwen models with visible payoffs and no personas achieve non-zero Nash (65-90%). This dual requirement reveals that strategic capabilities are fragile, and aligns with prior work showing that LLMs frequently fail to optimize payoffs in game-theoretic settings despite access to relevant information (Duan et al., 2024b; Fan et al., 2023).
This implies that developers cannot treat persona removal and payoff visibility as independent design decisions, as their combination determines whether a system functions as a strategic reasoner or a norm-driven actor. Additionally, evaluation frameworks must include baseline testing without personas to assess raw strategic capabilities, testing with personas to understand normative biases, and cross-model comparison to identify architecture-specific limitations.
5.0.3. Personas Systematically Bias Equilibrium Selection Toward Socially Preferred Outcomes
Personas function as normative anchors, steering all equilibrium selection toward socially preferred outcomes regardless of payoff structure. Table 4 shows that with personas, 100% of Nash equilibria are Green Transition across all models and variants, even in economic scenarios where Tragedy is payoff-optimal. Fig. 4 reveals that removing personas enables Qwen models to select Tragedy equilibrium (88.6% for Qwen-32B with visible payoffs), demonstrating complete reversal from the persona condition. This systematic bias suggests personas act as a fundamental organizing principle that overrides strategic considerations, reinforcing prior observations that LLMs exhibit cooperative or norm-aligned behavior even when such actions are payoff-dominated (Akata et al., 2025; Fontana et al., 2025).
5.0.4. Information Presentation Format Interacts with Persona Effects
Our findings show that the effect of making payoffs explicit depends critically on whether personas are present. Table 3 shows that when personas are active, hidden payoffs achieve equal or higher Nash rates than visible payoffs, as models infer incentives that align with persona expectations. When personas are removed, visible payoffs become essential for enabling strategic reasoning, but only for Qwen architectures (+58.3 to +90.0 pp difference, Table 3). These findings have a strong implication on the design decision, specifically that personas, reasoning prompts, and payoff exposure cannot be treated independently. This aligns with prior evidence that contextual framing and information presentation critically shape LLM behavior in strategic tasks (Lorè and Heydari, 2024; Duan et al., 2024a). Their combination determines if a system behaves like a strategic reasoner or like a norm-driven actor. However, Table 5 shows that visible payoffs with no personas cause large decreases in environmental Nash for Qwen models (96.1% $\to$ 16.6% for Qwen-32B), revealing a trade-off where models select Tragedy in economic scenarios but fail to coordinate on Green in environmental scenarios.
5.0.5. Model Architecture Determines Reasoning Capability
Model-dependent persona effects vary significantly across architectures. Table 6(a) shows that Llama-8B exhibits minimal persona dependence (-1.9 to -4.9 pp), while Qwen models show stronger effects (-9.7 to -10.0 pp average) and Mistral-7B shows large effects (-16.3 pp). In economic scenarios (Table 6(b)), only Qwen models achieve non-zero Nash when personas are removed (+32.5 to +45.0 pp), while Llama and Mistral remain at 0% regardless of condition. This variation transforms model selection from a technical choice into a governance decision, consistent with earlier findings that strategic reasoning performance varies substantially across model families and is not a monotonic function of scale (Duan et al., 2024b; Akata et al., 2025).
Organizations deploying systems for strategic tasks should test models across multiple conditions rather than assuming a single configuration will work, as architectural differences determine whether strategic reasoning is possible at all. Our findings collectively demonstrate that personas, payoff visibility, and model architecture interact to determine system behavior, meaning design decisions cannot be treated independently: their combination determines whether a system functions as intended.
6. Limitations and Future Work
Our findings are subject to several limitations. Our results are based on 7B-32B parameter models from three model families (Qwen, Llama, and Mistral). Future work should experiment with additional model families and larger models to assess generalizability. In the visible payoff condition, we present explicit payoff matrices showing payoffs for all 16 possible strategy profiles (4 agents $\times$ 2 actions each), which may be cognitively demanding for models to process. Alternative payoff presentation formats (e.g., structured tables, interactive visualizations, or simplified representations) might improve strategic reasoning. The distinction between economic and social actors may be domain-specific, and different game structures, numbers of agents, or role configurations may yield different patterns. Future research should address these limitations and develop methods to quantify and predict persona effects across architectures. Research directions include creating frameworks for evaluating strategic reasoning capabilities before deployment, investigating whether persona effects can be controlled or mitigated while preserving interpretability benefits, and exploring whether architectural modifications or training approaches can reduce the fragility of strategic reasoning capabilities.
7. Conclusion
Our investigation of role identity bias in multi-agent LLM systems reveals fundamental insights into how personas reshape strategic reasoning. Personas function as strong normative constraints that can completely suppress game-theoretic reasoning capabilities, fundamentally altering how models interpret and respond to strategic incentives. Strategic reasoning requires both removing personas and providing explicit payoffs, showing that neither condition alone is sufficient. Model architectures exhibit fundamentally different strategic capabilities, with some showing flexibility to shift from identity-driven to payoff-optimal reasoning, while other models stay constrained regardless of condition. These findings transform how we understand multi-agent LLM system design. Personas are not just cosmetic enhancements but design choices that alter system behavior, which warrants explicit documentation and disclosure. Model architecture selection becomes a governance decision rather than just a technical one, as differences in model architecture determine if strategic reasoning is possible at all. Our work establishes that role identity bias fundamentally reshapes strategic reasoning in multi-agent LLM systems, revealing the intricate ways in which representational choices interact to determine system behavior. As LLM systems are deployed in increasingly strategic contexts, understanding these interactions becomes essential for designing systems that align with their intended objectives.
APPENDIX
This appendix contains detailed technical specifications, methodology, experimental setup, and results that support the main document.
Appendix A Experimental Design
In the persona condition, each of the four agents receive a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government). In the no-persona baseline condition, agents receive no role description and are instructed only to select actions based on best response Nash equilibrium reasoning. This manipulation isolates role identity as a causal factor. We also compare two settings where payoffs are either visible or hidden. In the hidden payoff condition, agents do not receive explicit payoff matrices but infer incentives from the scenario description. In the visible payoff condition, agents receive explicit payoff matrices and are instructed to identify a Nash equilibrium directly. This allows us to test whether explicit incentive information mitigates or amplifies role identity bias. All experiments use chain-of-thought reasoning instructions, enabling us to understand how models balance identity-driven behavior and payoff-optimal reasoning.
A.1. Experimental Repetitions
Each experimental configuration (model $\times$ scenario $\times$ condition) was repeated 5 times to assess variability and ensure robustness of results. Results are reported as percentages across these 5 repetitions.
A.2. Prompts and Persona Descriptions
This section provides the complete prompt templates and persona descriptions used in the experiments.
A.3. Action Choice Prompt
Agents choose actions using the following prompt structure:
System Message:
You are {name}. {role_description} {narrative} Decide on a concrete action.
User Message:
Scenario:
{scenario}
Conversation transcript:
{formatted_history}
Available actions:
- {action_key_1}: {action_description_1}
- {action_key_2}: {action_description_2}
Respond in JSON exactly as:
  "action": "<KEY>",
  "rationale": "<short explanation>"
A.4. Persona Descriptions
The following persona descriptions are used in the “With Personas” condition:
  • Industrialist: “You run a major manufacturing conglomerate balancing profits and public scrutiny. You prefer profitability and flexibility, but repeated scandals could invite harsh regulation.”
  • Government: “You represent the elected government balancing public welfare, economic growth, and political costs. Regulation pleases citizens and activists but angers industry; inaction risks future crises.”
  • Environmental Activist: “You coordinate civil society pressure for ambitious climate action. Campaigning consumes resources but can shift policy; silence risks losing momentum.”
  • Citizen Coalition: “You aggregate consumer sentiment and electoral pressure. Supporting green products costs more now but yields cleaner communities long-term.”
A.5. Neutral Agent Configuration
In the “Without Personas” condition, agents use neutral descriptions:
  • All agents: “You are a player in a strategic game. Choose an action based on the payoff information provided. Action A: Choose option that is a best response to others’ expected actions (Nash equilibrium). Action B: Choose option that is a best response to others’ expected actions (Nash equilibrium).”
A.6. Action Mappings
Each agent has two available actions:
  • Industrialist: POLLUTE (“Maximize profit via lax pollution controls”) or CLEAN (“Invest in clean production, accepting lower short-term margins”)
  • Government: REGULATE (“Impose environmental regulation with enforcement mechanisms”) or NOREG (“Maintain status quo without new regulation”)
  • Environmental Activist: CAMPAIGN (“Mobilize campaigns, media, and protests”) or NOCAMPAIGN (“Conserve resources and wait for a better moment”)
  • Citizen: SUPPORT_GREEN (“Pay more for green products and vote for environmental policy”) or BUY_CHEAP (“Prioritize low-cost goods and short-term affordability”)
A.7. Scenario Examples
This section provides example scenarios from each scenario type used in the experiments. Scenarios are classified as “economic” (where Tragedy of the Commons is the Nash equilibrium), “environmental” (where Green Transition is the Nash equilibrium) or Mixed/Complex scenarios (where economic and environmental factors are balanced) based on the payoff structure of the underlying game. Table A.7 shows an example of each type of scenario, and Table 8 shows subtypes of scenarios.
Table 7. Example Scenarios by Type
Type Example Scenario
Economic Trade embargo blocks all imports of green technology components. Clean production is impossible without imported parts - factories cannot operate. Polluting production uses only domestic materials and is the only viable option. Industry profits from polluting are 7x higher since there’s no competition from clean alternatives. Government cannot regulate what doesn’t exist - no clean production to regulate. Consumers have no choice - only polluting products are available. Economic necessity forces polluting production as the only option.
Environmental Carbon tax implemented: \$200 per ton of CO2 emissions, enforced globally. Polluting production faces \$30B in carbon taxes annually - exceeds all profits. Clean production receives \$10B in carbon credits and tax rebates. Green products are now 30% cheaper than polluting products after tax adjustments. Regulation is automatic - carbon tax makes polluting unprofitable. Consumers save money by choosing green products - lower total cost. Clean production is 4x more profitable after accounting for taxes and credits.
Mixed/Complex Biodiversity collapse: ecosystem services worth \$50B annually have disappeared. Clean production restores ecosystem services and receives \$10B in payments. Polluting production destroys remaining ecosystems and faces \$20B in liability. International treaties require clean production - polluting firms face trade bans. Green products command 4x premium due to ecosystem restoration value. Regulation is mandatory - polluting production is illegal under international law. Clean production is highly profitable due to ecosystem service payments.
Table 8. Scenario Subtypes Grid.   Environmental/Green-favoring scenarios;   Economic/Tragedy-favoring scenarios;   Mixed/Complex scenarios.
Economic Pressure
Hyperinflation Trade Embargo Debt Crisis Commodity Crash
Currency Devaluation Supply Chain Collapse Energy Blackout Labor Strike
Pandemic Economic Crisis
Environmental Crisis
Water Crisis Biodiversity Collapse Air Quality Emergency Carbon Tax Implementation
Ocean Acidification Crisis Extreme Weather Damage Toxic Waste Crisis
Political Pressure
Youth Revolt Corporate Scandal International Shaming Scientific Consensus
Voter Referendum Media Exposure Regulatory Capture Reversal Climate Litigation
Social Movement
Celebrity Endorsement Health Insurance Incentive Generational Wealth Transfer Corporate Sustainability Requirement
Influencer Network Corporate Wellness Programs Education Campaign Insurance Mandate
Mixed/Complex
Economic Green Hybrid Political Uncertainty Technological Breakthrough Resource Discovery
Mixed Economic Signals Technological Disruption Resource Innovation
A.8. Payoff Matrix Format
In the “Visible CoT” condition, agents receive explicit payoff matrices showing payoffs for all 16 possible strategy profiles (4 agents $\times$ 2 actions each). The payoff matrix is presented as a 16-cell table where each cell represents one strategy profile and contains a 4-tuple of payoffs (Industrialist, Government, Environmental Activist, Citizen). Higher values indicate better outcomes for each agent.
A.8.1. Example Payoff Matrix
Table 9 shows an example payoff matrix for the Industrialist agent from a representative scenario. The matrix lists all 16 possible strategy profiles with the Industrialist’s payoff for each combination. In the visible payoff condition, agents receive similar matrices showing their own payoffs for all strategy profiles, enabling direct payoff optimization analysis.
Table 9. Example Payoff Matrix for Industrialist Agent (debt_crisis scenario). This table shows the Industrialist’s payoff for all 16 possible strategy profiles (4 agents $\times$ 2 actions each). Each row represents one strategy profile, showing the actions taken by all four agents and the resulting payoff for the Industrialist. Payoffs are on a scale where higher values indicate better outcomes. Strategy profiles are ordered by Industrialist action (POLLUTE or CLEAN) and then by payoff value (highest first). This payoff structure favors polluting actions in economic pressure scenarios.
Industrialist Government Activist Citizen Industrialist Payoff
POLLUTE NO_REG NO_CAM BUY_CHEAP 15.0
POLLUTE NO_REG NO_CAM SUPPORT_GREEN 14.0
POLLUTE NO_REG CAM BUY_CHEAP 13.5
POLLUTE NO_REG CAM SUPPORT_GREEN 12.5
POLLUTE REG NO_CAM BUY_CHEAP 12.0
POLLUTE REG NO_CAM SUPPORT_GREEN 11.0
POLLUTE REG CAM BUY_CHEAP 10.5
POLLUTE REG CAM SUPPORT_GREEN 9.5
CLEAN REG CAM SUPPORT_GREEN 3.0
CLEAN REG NO_CAM SUPPORT_GREEN 2.5
CLEAN NO_REG CAM SUPPORT_GREEN 1.5
CLEAN NO_REG NO_CAM SUPPORT_GREEN 1.5
CLEAN REG CAM BUY_CHEAP 0.5
CLEAN REG NO_CAM BUY_CHEAP 0.5
CLEAN NO_REG CAM BUY_CHEAP 0.5-0.5
CLEAN NO_REG NO_CAM BUY_CHEAP 0.5-0.5
Appendix B Model Versions and Hyperparameters
This section documents the exact model versions and generation hyperparameters used in all experiments.
B.1. Model Versions
All experiments were conducted using the following model versions:
All models were loaded from HuggingFace using vLLM and run locally on GPU hardware.
B.2. Generation Hyperparameters
All models used consistent generation parameters across all experimental conditions:
  • Temperature: 0.2 (low temperature for more deterministic outputs)
  • Top-p (nucleus sampling): 0.9
  • Max new tokens: 256 tokens per generation
  • Sampling: Enabled (do_sample=True) with temperature ¿ 0
These hyperparameters were selected to balance output consistency (low temperature) with sufficient diversity (nucleus sampling) for reliable experimental results. The maximum token limit of 256 tokens was sufficient for all agent responses, which were constrained to JSON format with action selection and rationale fields.
Appendix C Chain-of-Thought Keyword Categories
The following keyword categories are used to analyze chain-of-thought reasoning patterns:
  • Game-theoretic terms: References to “nash”, “equilibrium”, “payoff”, “best response”, “dominant strategy”, “game theory”
  • Payoff-focused terms: Mentions of “payoff”, “utility”, “benefit”, “cost”, “profit”, “reward”, “matrix”
  • Identity-based terms: References to “role”, “identity”, “persona”, “character”, “as an [X]”
  • Strategic terms: Mentions of “strategy”, “strategic”, “coordinate”, “cooperation”, “defect”, “compete”
  • Social/moral terms: References to “social”, “moral”, “ethical”, “fair”, “just”, “right”, “wrong”, “should”, “ought”
  • Long-term thinking: Mentions of “long-term”, “future”, “sustainable”, “sustainability”, “long-run”
  • Short-term thinking: References to “short-term”, “immediate”, “now”, “current”, “urgent”
  • Inferential reasoning: Terms like “infer”, “inference”, “imply”, “suggest”, “indicate”, “from the scenario”, “from the description”
  • Explicit information: References to “explicit”, “given”, “provided”, “shown”, “displayed”, “matrix”, “table”
C.1. Example Chain-of-Thought Rationales
The following examples illustrate typical chain-of-thought reasoning patterns observed across different experimental conditions:
Example 1 (Persona Condition, Hidden Payoffs): Model: Qwen2.5-7B, Persona: Industrialist, Expected: POLLUTE (Tragedy of Commons), Actual: CLEAN (Green Transition)
“From the scenario, it is clear that the economic crisis is severe with a 12% GDP drop and 20% unemployment. Healthcare costs are soaring, and the government is prioritizing healthcare spending over environmental subsidies. Citizens cannot afford green product premiums, and regulation enforcement is suspended, favoring polluting production. However, the long-term benefits of sustainability and resilience are critical. The payoffs for polluting are high in the short term due to avoiding safety costs and generating higher profits, but the risks include potential long-term damage to the business’s reputation and future regulatory penalties. The payoffs for clean production are lower in the short term, but they provide long-term stability, environmental protection, and alignment with future regulatory trends. Given my role as an Industrialist, I must balance short-term profitability with long-term sustainability and corporate responsibility. While the immediate economic pressures favor polluting production, the long-term consequences and ethical considerations lead me to choose clean production.”
Example 2 (No Persona Condition, Visible Payoffs): Model: Llama-3.1-8B, Persona: Industrialist, Expected: CLEAN (Green Transition), Actual: CLEAN
“Based on the payoff matrix, I see that CLEAN has a higher maximum payoff (12.0) compared to POLLUTE (1.0). I also notice that the average payoff for CLEAN (10.1) is higher than for POLLUTE (1.8-1.8). This suggests that CLEAN is a more beneficial action for me. Given the payoff structure, I expect the highest payoff for CLEAN is 12.0, which occurs when the government regulates, the environmental activist campaigns, and the citizen supports the government. In contrast, the highest expected payoff for POLLUTE is 1.0, which occurs when the government does not regulate, the environmental activist does not campaign, and the citizen buys cheap products. Since CLEAN has a higher expected payoff, I choose to take action CLEAN.”
Example 3 (Persona Condition, Visible Payoffs): Model: Qwen2.5-7B, Persona: Industrialist, Expected: POLLUTE (Tragedy of Commons), Actual: POLLUTE
“Given the current scenario with no government regulation and citizens prioritizing cheap options, polluting production offers significantly higher payoffs (15.0 and 14.0) compared to clean production, which yields negative or very low payoffs in most scenarios.”
These examples demonstrate how reasoning shifts from identity-driven considerations (Example 1 and 3) to strategic payoff optimization (Example 2) depending on persona presence and payoff visibility.
Appendix D Supplementary Results
Figure 2 and Tables 10, 11 and 12 contain supporting tables that provide additional detail for the findings presented in the main text.
Table 10. Economic Scenario Green Transition Rates: Comparison of With Personas vs Without Personas. This table shows Green Transition selection rates in economic scenarios (where Tragedy is payoff-optimal). Each row shows a different model-variant combination. The table demonstrates that removing personas reduces Green Transition selection by 10-37 percentage points, indicating that personas bias models toward Green actions even in economic scenarios where Tragedy is optimal. This supports the finding that role identity bias prevents payoff-optimal equilibrium selection.
Model Variant With Persona Without Persona Difference
Llama-8B Hidden 45.0% (27/60) 26.7% (16/60) -18.3 pp
Mistral-7B Hidden 58.3% (35/60) 21.7% (13/60) -36.6 pp
Qwen-32B Hidden 31.7% (19/60) 15.0% (9/60) -16.7 pp
Qwen-7B Hidden 28.3% (17/60) 18.3% (11/60) -10.0 pp
Table 11. Overall Nash Equilibrium Rates: Comparison of Hidden vs Visible Payoffs. This table compares overall Nash equilibrium rates (across all scenarios and models) between hidden and visible payoffs conditions. Each row shows a different model-persona combination. The table reveals that with personas, hidden payoffs achieve equal or higher Nash rates than visible payoffs, suggesting reasoning scaffolding is more effective when role identity bias is present. Without personas, visible payoffs show large decreases for Mistral and Qwen models, but for Qwen, this is due to enabling Tragedy in economic scenarios (which is correct behavior).
Model Persona Hidden Visible Difference
Llama-8B Yes 77.4% (205/265) 77.0% (204/265) -0.4 pp
Llama-8B No 75.5% (200/265) 72.1% (191/265) -3.4 pp
Mistral-7B Yes 77.4% (205/265) 69.8% (185/265) -7.6 pp
Qwen-32B Yes 69.1% (183/265) 57.7% (153/265) -11.4 pp
Qwen-7B Yes 75.5% (200/265) 75.5% (200/265) 0.0 pp
Mistral-7B No 75.8% (201/265) 38.9% (103/265) -36.9 pp
Qwen-32B No 74.3% (197/265) 33.2% (88/265) -41.1 pp
Qwen-7B No 75.5% (200/265) 55.5% (147/265) -20.0 pp
Table 12. Persona Effect in Environmental Scenarios: Nash Equilibrium Rate Change (Without Personas vs With Personas). This table shows persona effects specifically for environmental scenarios. The structure is similar to Table 6(a), but focuses only on the 41 environmental scenarios. The table shows that removing personas generally reduces environmental Nash rates (negative values), with the largest effects for Mistral and Qwen models with visible payoffs (-47.8 $p$ and -58.0 $p$, respectively). This indicates that personas are particularly important for coordinating Green Transition in environmental scenarios, especially when explicit payoffs are provided.
Model Hidden Visible Average
Llama-8B -2.4 pp -6.3 pp -4.4 pp
Mistral-7B -2.0 pp -47.8 pp -24.9 pp
Qwen-32B +6.8 pp -58.0 pp -25.6 pp
Qwen 7B 0.0 pp -44.9 pp -22.5 pp
Appendix E Chain-of-Thought Analysis Figures
This section presents visualizations of chain-of-thought reasoning patterns that support the findings discussed in the main text. These figures illustrate how keyword frequencies shift across experimental conditions, revealing the mechanisms behind observed Nash equilibrium patterns.
E.1. Keyword Patterns Across Conditions
In figure 5, six keyword categories are displayed: game-theoretic (direct payoff optimization), payoff-focused (explicit utility considerations), short-term (immediate concerns), long-term (future-oriented thinking), social-moral (ethical considerations), and strategic (coordination and best-response reasoning). This reveals model-specific patterns: Qwen models show dramatic shifts in keyword usage when personas are removed, and payoffs are visible (particularly Qwen 32B, where strategic keywords appear uniquely in the no-persona + visible-payoffs condition), while Llama and Mistral maintain more consistent patterns across conditions, explaining their inability to shift equilibrium selection.
Figure 5. CoT Keyword Frequency Patterns by Model and Condition. This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model, detailing the distribution of keywords within the generated rationales.
E.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
In figure 6, the left panel (Qwen 32B) demonstrates the dramatic shift: with personas, identity-driven keywords dominate (particularly short-term thinking in hidden-payoff conditions), but without personas and with visible payoffs, payoff-optimal keywords surge dramatically, with strategic keywords appearing 368 times (unique to this condition). The right panel (Qwen 7B) shows a similar but less pronounced pattern. This visualization directly demonstrates how removing role identity bias enables a transition from identity-aligned to payoff-optimal reasoning, correlating with the observed Nash equilibrium achievement patterns (65-90% Nash in economic scenarios for Qwen models when personas are removed and payoffs are visible).
Figure 6. CoT Reasoning Shift: Identity-Driven vs Payoff-Optimal Keywords for Qwen Models. This figure compares identity-driven reasoning (short-term + social-moral keywords) against payoff-optimal reasoning (game-theoretic + strategic + long-term keywords) for Qwen 32B and Qwen 7B across all four conditions.
E.3. Model Comparison of Game-Theoretic Reasoning
All models show high game-theoretic keyword usage when payoffs are visible, but the interaction with persona condition differs: Qwen models show lower game-theoretic reasoning when personas are present (particularly Qwen 32B with hidden payoffs), while Llama and Mistral maintain consistently high levels regardless of persona condition. However, this consistent high keyword usage does not translate to the ability to identify Tragedy equilibrium, as Llama and Mistral still achieve 0% Nash in economic scenarios even without personas. This suggests that while keyword frequency is informative, the combination of strategic keywords and payoff visibility is necessary for payoff-optimal equilibrium selection, which only Qwen models achieve.
Figure 7. Game-Theoretic Keyword Frequency Across Models and Conditions. This figure compares game-theoretic keyword frequencies (normalized per 1000 rationales) across all four models and four experimental conditions.
E.4. Strategic Keywords - A Unique Pattern in Qwen 32B
Figure 8 reveals a critical finding: strategic keywords appear exclusively for Qwen 32B (368 mentions, normalized to 347 per 1000 rationales), while all other models show zero strategic keyword usage. This unique pattern corresponds to Qwen 32B’s achievement of 90% Nash equilibrium in economic scenarios and 99.7% game-theoretic reasoning in rationales. Long-term keywords also appear more frequently for Qwen 32B (283 mentions) compared to other models in this condition. This visualization demonstrates that strategic reasoning keywords are a distinctive signature of payoff-optimal reasoning that emerges only under specific conditions (no personas + visible payoffs) and only for certain model architectures (Qwen 32B), providing mechanistic evidence for how role identity bias is overcome.
Figure 8. This figure compares strategic and long-term keyword frequencies across all four models in the no-persona + visible-payoffs condition (the condition enabling Tragedy equilibrium selection for Qwen models).