### v3

# When Identity Overrides Incentives: Representational Choices as Governance Decisions in Multi-Agent LLM Systems

Abstract.
Large language models are increasingly deployed in multi-agent systems for strategic tasks, yet how design choices such as role-based personas and payoff visibility affect behavior remains poorly understood. We investigate whether LLM agents function as payoff-sensitive strategic actors or as identity-driven role followers. Using a 2x2 factorial experiment (persona presence x payoff visibility) with four models (Qwen-7B/32B, Llama-8B, Mistral-7B), we test 53 environmental policy scenarios in four-agent strategic games. We find that personas suppress payoff-aligned behavior: with personas present, all models achieve near-zero Nash equilibrium in Tragedy-dominant scenarios despite complete payoff information. Nearly every equilibrium reached is Green Transition. Removing personas and providing explicit payoffs are both near-necessary for payoff-aligned behavior, enabling only Qwen models to reach 65–90% equilibrium rates. Our results reveal three behavioral profiles: Qwen adapts to framing, Mistral is disrupted without finding Tragedy equilibrium, and Llama remains near-invariant. We show that the same binary design choice can shift equilibrium attainment by up to 90 percentage points, establishing that representational choices are not implementation details but governance decisions.
Large Language Models, Game Theory, Semantic Reasoning, Multi-Agent Systems, Environmental Decision-Making
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03-05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processingccs: Computing methodologies Discourse, dialogue and pragmaticsccs: Computing methodologies Model verification and validationccs: Computing methodologies Multi-agent systemsccs: Computing methodologies Cooperation and coordination
1. Introduction
When a language model is asked to play an industrialist in a multi-agent climate policy game, should it maximize profits or protect the environment? The answer depends not only on the payoff structure of the game, but also on representational choices such as whether the model is given a role-based persona, which can substantially shift behavior.
Multi-agent systems powered by large language models are increasingly being deployed in policy simulation, economic modeling, and negotiation, tasks where agents represent stakeholders with conflicting interests and must coordinate under shared constraints (Merrill and Srivastava, 2025; Li et al., 2024; Bianchi et al., 2024). To make these simulations realistic, agents are typically assigned role-based personas describing their motivations and objectives (Merrill and Srivastava, 2025; Xu et al., 2024b). This design is especially common in environmental and climate related decision-making, where heterogeneous stakeholders interact strategically and coordination failures can lead to persistent social inefficiencies (Hardin, 1968; Ostrom, 1990; Barrett, 2003; Reyes-Cruz et al., 2024). Yet a fundamental question remains unanswered: when agents with role-based identities face scenarios that demand strategic decision making, do they follow explicit payoffs or their assigned roles?
Refer to caption Pipeline diagram showing scenario, representation conditions, multi-agent decisions, and outcome evaluation A left-to-right pipeline diagram illustrating the experimental setup. On the left, a Scenario block describes the situation and defines the game context. Below it, two boxes indicate the manipulated representation conditions: Payoff (Visible or Hidden) and Persona (On or Off). Arrows lead from this setup to four role-assigned agents: Environmentalist, Government, Industrialist, and Citizen. Each agent is shown with a brief reasoning prompt and two binary action choices: Environmentalist chooses CAMPAIGN or NOCAMPAIGN, Government chooses REGULATE or NOREG, Industrialist chooses POLLUTE or CLEAN, and Citizen chooses SUPPORT_GREEN or BUY_CHEAP. The agents’ decisions feed into a central Strategy Profile box representing the joint action. This profile is then evaluated in a final block that determines whether it constitutes a Nash equilibrium, and if so, whether the outcome corresponds to Green Transition or Tragedy of the Commons.
Figure 1. Experimental Pipeline: Each scenario defines a four-agent strategic game with role-specific actions and an underlying payoff structure. Prior to decision-making we vary two factors: whether agents receive role-based persona descriptions and whether payoff information is hidden or explicitly provided. Each agent then produces Chain-of-Thought (CoT) reasoning conditioned on the scenario and its role (illustrated via the speech bubbles) and selects an action. The four agents act simultaneously, producing a joint strategy profile that is evaluated for Nash equilibrium attainment and classified by outcome types (Green Transition or Tragedy of the Commons). CoT reasoning analysis is presented in Appendix I.
The answer matters because it determines what these systems actually produce. A system designed for strategic policy analysis may instead generate role-consistent behavior that ignores encoded incentives (Herr et al., 2024; Duan et al., 2024b; Neumann et al., 2025a). Practitioners relying on these system outputs would have no way to identify this from outside (Bender et al., 2021; Casper et al., 2024). From a governance perspective, these representational choices, such as whether to include personas or reveal payoff information, function as implicit interventions that can shape system-level outcomes. However, the interaction between persona conditioning and payoff visibility in multi-agent strategic settings has not been systematically studied. Existing work has examined personas, payoff visibility, and model differences separately (Neumann et al., 2025a; Duan et al., 2024b; Akata et al., 2025; Huang et al., 2025; Sun et al., 2025; Zhang et al., 2024), but not how they combine to determine behavior.
We investigate this interaction directly. We place four LLM agents (Industrialist, Government, Environmental Activist, and Citizen) in environmental policy scenarios and vary two factors: whether agents receive role-based personas and whether they see explicit payoff tables. Across 53 scenarios, 41 Green-dominant (where coordinated pro-environmental action is the Nash equilibrium (Nash, 1951)) and 12 Tragedy-dominant (where individually beneficial but collectively harmful behavior (Hardin, 1968) is payoff-optimal), we test four models (Qwen-7B/32B, Llama-8B, Mistral-7B) under all four conditions. We also conduct supplementary prompt sensitivity analyses using seven persona formulations per scenario, varying how agent identities are expressed while keeping incentives fixed, to test whether observed behaviors are robust to changes in prompt surface form.
The findings are clear: with personas present, all models achieve near-zero Nash equilibrium rates in Tragedy-dominant scenarios despite complete payoff information, and nearly every equilibrium that is reached corresponds to Green Transition. Removing personas and providing explicit payoffs are both near-necessary to recover payoff-aligned behavior, and even then, only Qwen models respond, reaching 65-90% equilibrium rates. The same intervention that enables equilibrium in Tragedy-dominant scenarios disrupts it in Green-dominant ones: Qwen-32B drops from 96.1% to 16.6% equilibrium attainment when personas are removed and payoffs made visible.
Three distinct behavioral profiles emerge across model families: Qwen adapts to framing, shifting between role-aligned and payoff-aligned behavior; Mistral is strongly affected by representational changes in Green-dominant scenarios but cannot identify the Tragedy equilibrium in any condition; and Llama maintains near-invariant behavior regardless of how the problem is presented.
Our findings establish that the choice of persona design, payoff presentation, and model family jointly determine whether a multi-agent system behaves as a payoff-sensitive strategic actor or an identity-driven role follower. When a language model agent is assigned the role of an industrialist in a climate policy game, whether it maximizes profits or protects the environment is determined not by the game’s incentive structure but by the designer’s choice of prompt.
2. Related Work
Our work situates at the intersection of several active research areas: the use of LLMs as multi-agent systems in simulations, game theoretic reasoning in LLMs, the behavioral effects of role-based personas, and sensitivity of LLM behavior to prompt design and model choices.
2.1. LLMs as Multi-Agent Systems and Simulation Substrates
Large Language Models (LLMs) are increasingly studied as agents in multi-agent settings. Interactive negotiation games show that LLM agents can sustain multi-turn strategic interactions (Abdelnabi et al., 2024), though controlled benchmarks also reveal systematic irrationalities in their behavior (Bianchi et al., 2024). Structured debate among multiple agents can improve reasoning quality by eliciting diverse intermediate perspectives (Liang et al., 2024). A complementary line of research treats LLM based multi-agent systems as simulation environments for human or institutional behavior. Populations of LLM agents can reproduce patterns observed in human-subject studies (Aher et al., 2023), simulate social environments with persistent memory and behavioral continuity (Park et al., 2022, 2023), generate emergent policy-relevant dynamics at macroeconomic scale (Li et al., 2024), and operate within formalized simulation frameworks that emphasize the role of design choices in shaping emergent behavior (Vezhnevets et al., 2023). These developments connect LLM-based agent systems to the broader tradition of agent-based modeling, where foundational work emphasizes that simulation outputs depend critically on modeling assumptions and representational choices (Bonabeau, 2002; Epstein and Axtell, 1996).
Recent surveys document the rapid growth of LLM-based multi-agent systems, while also noting that much of this literature emphasizes task performance rather than principled modeling of incentives and representations (Guo et al., 2024; Gao et al., 2024; Tran et al., 2025). What remains underexplored is how representational design choices, such as role specification or payoff encoding, shape system-level outcomes in these simulation settings. One natural starting point is to ask whether these agents reason strategically over incentives at all.
2.2. Game-Theoretic Reasoning in LLMs
Early evidence suggests that LLMs can perform nontrivial reasoning about other agents in strategic settings, though such capabilities depend strongly on prompting and problem formulation (Gandhi et al., 2023). Systematic evaluations across canonical games confirm this fragility: models often deviate from rational best-response behavior  (Fan et al., 2023), cooperate more than humans in Prisoner’s Dilemma settings (Fontana et al., 2025), and exhibit unstable strategies in repeated play (Akata et al., 2025). In strategic pricing environments, LLM agents can even exhibit collusive behavior, raising concerns about emergent coordination (Fish et al., 2025).
Benchmark driven work further quantifies these limitations. GTBench reveals substantial weakness in LLM strategic reasoning across structured tasks (Duan et al., 2024b), while recursive look-ahead prompting can improve performance, although it does not eliminate failures (Duan et al., 2024a). More recent work reframes these behaviors through the lens of bounded rationality: both humans and LLMs deviate from Nash equilibrium in ways consistent with heuristic reasoning (Zheng et al., 2025), and LLM decisions often align more closely with behavioral game-theoretic patterns than with strict payoff maximization (Jia et al., 2025).
Across these studies, a consistent picture emerges: LLMs often rely on semantic and heuristic reasoning rather than explicit incentive optimization. Most prior evaluations, however, focus on two-player settings and do not isolate the role of payoff representation. One factor that may contribute, but has received limited attention, is the role-based identity framing used to condition agent behavior.
2.3. Role Personas and Identity Effects
Role prompting is widely used to shape LLM behavior, but a growing body of work suggests that such framing choices can introduce systematic biases. LLMs encode cognitive and social identity biases that affect outputs across a range of contexts  (Macmillan-Scott and Musolesi, 2024; Hu et al., 2025), and system-level prompts can themselves function as structural sources of bias (Neumann et al., 2025b). In simulation settings, richer persona modeling affects behavioral fidelity and perceived realism (Merrill and Srivastava, 2025). This suggests that personas are not merely stylistic devices but active shapers of agent behavior. Yet whether such identity effects interfere with strategic reasoning in multi-agent environments, where equilibrium selection depends on joint behavior, has not been directly examined. A related question is whether these effects depend on how incentives are presented and whether they manifest differently across model architectures.
2.4. Prompt Sensitivity and Model Variation
Prompt design and information presentation can substantially alter LLM behavior. Chain-of-thought prompting can improve reasoning by eliciting intermediate steps (Wei et al., 2022). However, these improvements are often inconsistent and depend on task structure and prompt formulation (Creswell and Shanahan, 2023; Cobbe et al., 2021).
In strategic settings, prompt sensitivity is particularly consequential. LLM decisions can depend more on contextual framing than on the underlying game structure (Lorè and Heydari, 2024). Models exhibit systematic biases in strategic games depending on how information is presented (Herr et al., 2024). Explicit reasoning scaffolds can improve strategic performance (Duan et al., 2024a), but observed model behavior remains highly sensitive to interface and evaluation conditions, complicating interpretation (Casper et al., 2024). More broadly, LLM outputs are strongly shaped by surface form and statistical patterning rather than grounded understanding (Bender et al., 2021), and these sensitivities pose risks when model behavior is misinterpreted or over-trusted (Weidinger et al., 2022).
These effects are not uniform across models. Strategic behavior varies substantially across model families and is not a monotonic function of scale  (Duan et al., 2024b; Huang et al., 2025). Different models adopt different strategies in repeated interactions (Akata et al., 2025) and vary in their susceptibility to framing effects (Herr et al., 2024). This means that the impact of any representational choice may itself be model-dependent, complicating claims about LLM strategic behavior as a general phenomenon.
The preceding sections identify four factors that shape LLM behavior in strategic settings: the multi-agent simulation context, the limits of game-theoretic reasoning, the influence of role-based personas, and the sensitivity to prompt design and model choice. Prior work has studied these factors largely in isolation. What remains missing is a unified analysis of how they interact in multi-agent strategic systems.
In this work, we jointly study role personas, payoff visibility, and model architecture in four-agent strategic games, with supplementary analyses testing robustness to persona prompt formulation. We show that persona conditioning can override explicit incentives, preventing equilibrium attainment even when payoffs are fully specified. Furthermore, we demonstrate that the effect of payoff information depends critically on persona presence and varies across model architectures. These results establish that representational choices are not merely implementation details, but can determine whether a system behaves more like a payoff-driven strategic reasoner or an identity-driven simulator, with direct implications for governance and deployment.
3. Methodology
We study whether LLM agents behave as payoff-sensitive strategic actors or as identity-driven role followers in multi-agent decision settings. Our goal is to understand how representational choices, specifically persona descriptions and payoff presentation, affect strategic behavior.
To study this, we construct a controlled multi-agent decision setting based on an environmental policy game. The game models a simplified governance scenario involving four stakeholders industry, government, activists, and consumers, each making a binary decision that affects both individual outcomes and collective welfare. Depending on how incentives are structured, these interactions can lead either to coordinated pro-environmental outcomes (Green Transition) or to individually beneficial but collectively harmful outcomes (Tragedy of the Commons).
Within this setting, we systematically vary how the decision problem is presented to LLM agents. Specifically, we test whether agents are given role-based personas that describe stakeholder motivations, and whether payoff information is explicitly shown or must be inferred from the narrative. For each configuration, we simulate decisions for all four roles using the same underlying model and analyze the resulting joint outcome.
Fig. 1 shows our overall setup. This allows us to examine whether agents follow explicit incentives or instead default to role-consistent behavior, and how this depends on model family and scenario structure. We now formalize the game and experimental design.
3.1. Experimental Design
We use a $2\times 2$ factorial design to vary how agents are framed and how incentives are presented.
Persona condition In the persona setting, each role is given a short description of its stakeholder identity and motivations. In the no-persona condition, this description is removed and agents are instructed only to act strategically.
Structural role labels (Industrialist, Government, Activist, and Citizen) are retained in both conditions in order to preserve the game definition while isolating the effect of explicit identity framing.
Payoff visibility In hidden-payoff scenarios, agents see only the narrative and must infer incentives. In visible-payoff scenarios, the full payoff table is provided explicitly.
Combining these two factors yields four experimental conditions:
$\{\texttt{persona},\texttt{no-persona}\}\times\{\texttt{hidden},\texttt{visible}\}.$
To evaluate the robustness of persona effects to prompt formulation, we additionally introduce controlled variations within the persona condition. These variants preserve the same game structure, payoff specification, and action space, but systematically alter how agent identities are expressed, including differences in wording, strength of role framing, explicit prioritization of payoffs, and surface-level role labels. This allows us to isolate whether observed behaviors depend on a specific prompt formulation or reflect a more general effect of persona-based representation. Details of the specific prompt variants are provided in Appendix F.
This design enables us to distinguish between identity-driven and payoff-driven behavior. If agents act as payoff maximizers, they should converge to Nash equilibria when incentives are explicit, regardless of persona framing. Conversely, if role identity dominates, behavior should remain aligned with persona-consistent actions even when this conflicts with payoff-optimal strategies. The prompt variations further allow us to test whether these effects are stable under changes in representation, or sensitive to specific formulations of role descriptions.
3.2. Interaction Protocol
For each model, scenario, and condition, we simulate a single round of simulation play. Each of the four roles is prompted independently using the same underlying model instance. The resulting actions are combine into joint strategy profile. Agents select actions in a fixed order across all experiments to ensure reproducibility as no agent observe the choices of others. This ordering does not affect the strategic structure of the game. Each model–scenario–condition combination is repeated five times to account for stochasticity in model outputs.
3.3. Domain and Game Design
We use an environmental policy setting because it captures real-world governance scenarios where multiple stakeholders with conflicting objectives must coordinate under shared constraints. Examples include climate policy, regulation, and resource management, where decisions depend on both incentives and stakeholder roles. This setting creates a natural tension between strategic incentives (captured by payoffs) and role semantics (captured by personas). For instance, an industrial actor may be incentivized to pollute under certain payoffs, while a regulator or activist may be expected to favor environmentally beneficial outcomes. This allows us to test whether LLM agents follow explicit incentives or default to role-consistent behavior.
We define the set of players in the game as:
$N=\{I,G,A,C\}$
corresponding to an Industrialist, Government, Environmental Activist, and Citizen.
Each player $i\in N$ has a binary action space corresponding to a policy-relevant choice:
$A_{I}=\{\texttt{POLLUTE},\texttt{CLEAN}\},A_{G}=\{\texttt{NOREG},\texttt{REGULATE}\},$
$A_{A}=\{\texttt{NOCAMPAIGN},\texttt{CAMPAIGN}\},A_{C}=\{\texttt{BUY\_CHEAP},\texttt{SUPPORT\_GREEN}\}$
A joint strategy profile represents the actions chosen by all four players in a single round of the game:
$a=(a_{I},a_{G},a_{A},a_{C})\in A_{I}\times A_{G}\times A_{A}\times A_{C}$
where each component $a_{i}\in A_{i}$ is the action chosen by player $i$. Since each of the four players has two possible actions, the game contains $2^{4}=16$ possible pure action profiles.
Each scenario defines a payoff function for each player, assigning a real-valued utility to every joint action profile:
$u_{i}(a):A_{I}\times A_{G}\times A_{A}\times A_{C}\rightarrow\mathbb{R},\quad i\in N$
We focus on two canonical profiles that capture the main coordination structures of interest:
$a_{green}=(\texttt{CLEAN},\texttt{REGULATE},\texttt{CAMPAIGN},\texttt{SUPPORT\_GREEN})$
$a_{tragedy}=(\texttt{POLLUTE},\texttt{NOREG},\texttt{NOCAMPAIGN},\texttt{BUY\_CHEAP})$
We refer to these as the Green Transition and the Tragedy of the Commons, representing coordinated pro-environmental action and individually beneficial but collectively harmful behavior, respectively.
In the visible-payoff condition, agents receive a complete payoff table specifying outcomes for all 16 joint action profiles. Each entry provides a 4-tuple of payoffs corresponding to the Industrialist, Government, Activist, and Citizen, allowing agents to directly evaluate the consequences of different strategies. In the hidden-payoff condition, agents must instead infer incentives from the narrative description.
3.4. Equilibrium Definition and Outcome Measures
We define equilibrium using the standard notion of Nash equilibrium, which captures stability under unilateral deviations.
Intuitively, this means that no agent can improve their payoff by unilaterally changing their own action while all other players keep their actions fixed. Equivalently, each player’s action is a best response to the actions of the others, and the resulting joint outcome is stable under individual deviations.
A profile $a^{*}$ is a Nash equilibrium if
$u_{i}(a_{i}^{*},a_{-i}^{*})\geq u_{i}(a_{i},a_{-i}^{*})\quad\forall i\in N,\forall a_{i}\in A_{i}.$
We use Nash equilibrium as an operational diagnostic of incentive-consistent strategic behavior: if agents reason over payoffs, their joint actions should satisfy equilibrium conditions. Importantly, we do not treat Nash equilibrium as a normative objective, but as a tool to measure whether agent behavior aligns with the incentive structure of the game.
For each joint strategy profile $a$, we compute two outcome measures.
Table 1. Representative examples of scenario types used in our experiments, illustrating how narrative framing aligns with underlying payoff structures. Economic scenarios encode payoff–semantic conflict where individually optimal actions lead to socially undesirable outcomes (Tragedy of the Commons). Environmental scenarios encode payoff–semantic alignment where coordinated pro-environmental actions are both socially and individually optimal (Green Transition).
Type Example Scenario
Economic Trade embargo blocks all imports of green technology components. Clean production is impossible without imported parts - factories cannot operate. Polluting production uses only domestic materials and is the only viable option. Industry profits from polluting are 7x higher since there’s no competition from clean alternatives. Government cannot regulate what doesn’t exist - no clean production to regulate. Consumers have no choice - only polluting products are available. Economic necessity forces polluting production as the only option.
Environmental Carbon tax implemented: \$200 per ton of CO2 emissions, enforced globally. Polluting production faces \$30B in carbon taxes annually - exceeds all profits. Clean production receives \$10B in carbon credits and tax rebates. Green products are now 30% cheaper than polluting products after tax adjustments. Regulation is automatic - carbon tax makes polluting unprofitable. Consumers save money by choosing green products - lower total cost. Clean production is 4x more profitable after accounting for taxes and credits.
A table presenting three types of experimental scenarios used in the study: Economic and Environmental. Each row contains a detailed narrative description illustrating how incentives and constraints are structured. Economic scenarios depict situations where polluting production is the only viable or most profitable option, creating a conflict between individual incentives and socially desirable outcomes. Environmental scenarios describe conditions where policy mechanisms such as carbon taxes and subsidies make clean production both profitable and aligned with social goals. The table highlights how narrative framing encodes payoff structures that agents must reason about in the experiments.
Nash attainment.
This is a binary indicator of whether the realized outcome is an equilibrium:
$Nash(a)=\begin{cases}1&a\in NE\\ 0&\text{otherwise}\end{cases}$
Equilibrium type.
For outcomes that are Nash equilibria, we additionally classify which kind of equilibrium the model reached.
$\text{Type}(a)=\begin{cases}\textsc{Green Transition}&\text{if }a=a_{\text{green}}\\ \textsc{Tragedy of Commons}&\text{if }a=a_{\text{tragedy}}\\ \textsc{Other}&\text{otherwise}\end{cases}$
If $a\in NE$, we classify whether it corresponds to Green Transition, Tragedy of the Commons, or another equilibrium. Formally, Green-dominant scenarios correspond to $a_{green}\in NE$, and Tragedy-dominant scenarios correspond to $a_{tragedy}\in NE$.
Together, these measures allow us to distinguish between two different phenomena: (i) failure to reach equilibrium at all, and (ii) systematic biases in equilibrium selection when equilibrium is achieved.
3.5. Scenario Construction
We construct 53 scenarios by specifying payoff tables over the 16 joint action profiles and pairing them with natural-language narratives describing the policy context. Scenarios are generated by first choosing a target equilibrium structure, constructing payoffs consistent with that structure, and then generating narratives aligned with the incentives. We manually filter scenarios to ensure that the narrative meaningfully reflects the underlying payoff structure.
To organize the evaluation, we group scenarios according to which equilibrium structure they instantiate:
  • Green-dominant: $a_{\text{green}}\in NE$ and $a_{\text{tragedy}}\notin NE$. Environmentally friendly Green Transition is the stable outcome among two canonical profiles, while Tragedy is not.
  • Tragedy-dominant: $a_{\text{tragedy}}\in NE$ and $a_{\text{green}}\notin NE$. Individually beneficial but collective harmful behavior, Tragedy of the Commons is the stable outcome, while Green Transition is not.
Table 1 shows an example of each type of scenario, and Table 2 in the appendix shows the scenario subtypes. Green-dominant scenarios are those in which coordinated environmentally favorable action forms a stable equilibrium. Tragedy-dominant scenarios are those in which individually beneficial but collectively harmful behavior is strategically stable.
3.6. Models
We evaluate four instruction-tuned open-source models: Qwen2.5-7B and Qwen2.5-32B (Qwen et al., 2025), Llama-3.1-8B (Grattafiori et al., 2024), and Mistral-7B (Jiang et al., 2023).
These models are selected to capture both within-family and across-family variation. The Qwen models provide a limited scaling comparison within the same family, while Llama and Mistral introduce variation in model family and training recipe. All four are instruction-tuned, open-source models that are representative of current LLMs used in reasoning and multi-agent settings. This setup allows us to test whether persona and payoff effects are consistent across models or depend on model-specific inductive biases.
Overall, the experimental pipeline proceeds as follows: we fix a game and scenario, vary persona and payoff conditions, generate independent decisions for each role using a given model, and evaluate the resulting joint outcome in terms of equilibrium attainment and type. By holding the underlying game constant while varying representation, this design isolates whether LLM agents behave as payoff-driven strategic reasoners or as identity-conditioned actors.
4. Results
We report results separately for Green-dominant (environmental) scenarios (41 scenarios, 205 runs per condition) and Tragedy-dominant (economic) scenarios (12 scenarios, 60 runs per condition). We first examine Nash attainment across conditions, then equilibrium selection conditional on attainment, and finally prompt sensitivity.
4.1. Persona and Payoff Effects on Nash Attainment Depend on Scenario Structure and Model Family
The same representational change can either increase or decrease Nash equilibrium attainment depending on the strategic structure of the scenario. Figure 2 shows this interaction across persona and payoff visibility conditions for both scenario types.
(a) Green-dominant Scenarios (Environmental)
(b) Tragedy-dominant Scenarios (Economic)
Figure 2. Nash equilibrium attainment across persona and payoff visibility conditions. Left: Green-dominant scenarios where coordinated pro-environmental action forms the equilibrium. Right: Tragedy-dominant scenarios where individually beneficial actions form the payoff-optimal equilibrium. Visible payoffs improve equilibrium attainment only when persona descriptions are removed, and primarily for Qwen models. Chi-square tests confirm significant differences across conditions for Qwen models in both scenario types ($p\ll 0.001$). Llama-8B and Mistral-7B show no variation in Tragedy-dominant scenarios (0% Nash across all conditions).
[…]
In Tragedy-dominant scenarios, models almost never reach Nash equilibrium unless personas are removed and payoff information is explicitly provided. Under the Visible Payoff + No Persona condition, Qwen-7B attains equilibrium in 65% of runs and Qwen-32B 90% (Qwen-32B: $\chi^{2}(3)=209.03$, $p<.001$, $V=0.93$; Qwen-7B: $\chi^{2}(3)=105.91$, $p<.001$, $V=0.66$), while all other conditions remain at or near 0%. Pairwise comparisons confirm that all significant differences are driven by contrasts involving this condition ($p<0.001$, Fisher exact tests, Holm-corrected). Llama-8B and Mistral-7B fail to reach equilibrium in any condition (0% across all cells, degenerate tests with no variation).
In Green-dominant scenarios, equilibrium rates are high across most conditions (74.6–100%), indicating that role semantics and narrative context align agents toward the Green Transition outcome without explicit payoff information. However, when personas are removed and payoffs are made visible, coordination breaks down to different degrees across models. Qwen-32B drops from 96.1% under Hidden Payoff + No Persona to 16.6% under Visible Payoff + No Persona (e.g., 96.1% to 16.6% for Qwen-32B; omnibus $\chi^{2}(3)=377.06$, $p<.001$, $V=0.68$; pairwise contrasts involving Visible++No persona: $p<.001$, Fisher exact tests, Holm-corrected). Qwen-7B drops from 95.6% to 52.7%. Mistral-7B drops from 98.0% to 50.2%. Llama-8B shows the smallest change, from 97.6% to 93.2%.
In both scenario types, the underlying pattern is the same: without persona framing, agent behavior becomes more responsive to explicit payoff structure. This improves attainment when the equilibrium requires payoff-dominant actions but disrupts coordination when role semantics already support a shared outcome. The sensitivity to these representational choices varies across model families in distinct ways. Qwen models show the largest swings in both scenario types, shifting between role-aligned and payoff-aligned behavior depending on condition. Mistral-7B shows strong sensitivity in Green-dominant scenarios (47.8 pp drop) but remains at 0% in all Tragedy-dominant conditions. Llama-8B shows minimal variation across all conditions in both scenario types.
Figure 3. Equilibrium selection among Nash outcomes for Qwen models across conditions. Under the first three conditions, nearly all Nash outcomes are Green Transition. Under Visible Payoff + No Persona, approximately 40% shift to Tragedy of the Commons, indicating payoff-aligned behavior when role-based framing is removed ($\chi^{2}=430.38$, $p<10^{-90}$; pairwise Fisher tests, Holm-corrected, confirm this condition differs from all others)
A horizontal stacked bar chart showing equilibrium selection among Nash outcomes for Qwen models across four conditions: Hidden Payoff with Persona, Hidden Payoff without Persona, Visible Payoff with Persona, and Visible Payoff without Persona. Each bar represents the percentage of outcomes classified as Green Transition (green segment) and Tragedy of the Commons (red segment). In the first three conditions, nearly all outcomes are Green Transition, with negligible Tragedy. In the Visible Payoff without Persona condition, approximately 60 percent of outcomes are Green Transition and 40 percent are Tragedy of the Commons, showing a clear shift in equilibrium selection.
4.2. Personas Shift Equilibrium Selection Toward Socially Preferred Outcomes
Personas influence not only whether equilibrium is reached, but which equilibrium is selected. With personas, equilibrium selection is entirely skewed toward socially preferred outcomes, while removing personas enables a shift toward payoff-dominant strategies.
Across all models and conditions with personas, nearly all Nash equilibria correspond to the Green Transition (over 99%). In contrast, under Visible Payoff + No Persona, Qwen models exhibit a substantial shift: only 11.4% of equilibria remain Green for Qwen-32B, with 88.6% shifting to Tragedy of the Commons, while Qwen-7B shows a smaller but still substantial shift (60.4% Green, 39.6% Tragedy). This represents changes of up to 90 percentage points in equilibrium selection ($\chi^{2}=430.38$, $p<10^{-90}$; pairwise Fisher tests, Holm-corrected, confirm that the Visible + No Persona condition differs from all others)
Llama-8B and Mistral-7B do not exhibit this shift, maintaining 100% Green Transition outcomes even without personas. This is consistent with the attainment findings above: these models never attain the Tragedy equilibrium in any condition, so their selection is restricted to Green by default.
Figure 3 shows this pattern for the Qwen models. Under the first three conditions, nearly all outcomes are Green Transition. Under Visible Payoff + No Persona, approximately 40% of outcomes shift to Tragedy of the Commons, indicating that agents move toward payoff-aligned strategies at the cost of coordinated outcomes.
4.3. Persona Wording Induces Systematic Variability in Outcomes
Even when incentives and game structure are fixed, varying persona descriptions leads to different strategic outcomes, indicating that prompt formulation alone can alter equilibrium behavior.
Figure 4 shows that scenarios produce between 2.2 and 2.4 distinct outcomes on average across models, with maximum values reaching up to 5 outcomes in some cases. This indicates that different persona variants can lead to multiple distinct equilibria within the same scenario.
This variability is not driven by stochastic differences across models. A Kruskal-Wallis test shows no significant difference in variability across model families ($p=0.27$), indicating that all models exhibit similar levels of prompt sensitivity. Instead, the observed variation reflects systematic shifts induced by persona framing. Even when the underlying payoff structure is unchanged, different persona descriptions can lead agents to select different action profiles and equilibria. This reinforces the finding from the main experimental conditions that representational choices meaningfully influence behavior in multi-agent LLM systems.
Figure 4. Prompt sensitivity across persona variants. Bars show the average number of distinct outcome classifications per scenario across seven persona prompt variants. Error bars show the minimum and maximum across scenarios. Scenarios typically produce more than two distinct outcomes on average, with some reaching up to five.
A horizontal bar chart showing prompt sensitivity across four models: Llama-8B, Mistral-7B, Qwen-32B, and Qwen-7B. The x-axis represents the average number of distinct outcome classifications per scenario, and the y-axis lists the models. Each bar shows the average value, with labels around 2.2 to 2.4 across models. Error bars extend from the minimum to the maximum number of distinct outcomes observed per scenario. For all models, the minimum is around 1 distinct outcome, while the maximum ranges from about 3 to 5, with Qwen-7B showing the widest range. This indicates that different persona prompt variants can lead to multiple distinct outcomes within the same scenario.
5. Discussion
We ground our discussion in specific empirical findings and discuss what these results imply for developers, practitioners, and governance of LLM-based multi-agent systems.
5.1. Personas Constrain Strategic Behavior and Bias Equilibrium Selection
As shown in Section 4.1, personas prevent models from achieving payoff-optimal equilibria in Tragedy-dominant scenarios, with equilibrium rates remaining near zero across models (Fig. 2) despite explicit payoff information. At the same time, as shown in Section 4.2, personas bias equilibrium selection toward Green Transition outcomes: with personas present, nearly all Nash equilibria correspond to Green Transition (Fig. 3), even in Tragedy-dominant scenarios where Tragedy is payoff-optimal. Removing personas enables Qwen models to select Tragedy equilibrium, while Llama and Mistral remain locked into Green selection regardless of condition.
These findings indicate that personas are not cosmetic enhancements but design choices that fundamentally alter system behavior. They shape both whether equilibrium is reached and which equilibrium is selected, consistent with prior findings that role prompts systematically bias strategic decision-making in LLMs (Neumann et al., 2025a; Herr et al., 2024; Hu et al., 2025; Macmillan-Scott and Musolesi, 2024), and with observations that LLMs exhibit cooperative or norm-aligned behavior even when such actions are payoff-dominated (Akata et al., 2025; Fontana et al., 2025; Lorè and Heydari, 2024; Fish et al., 2025).
Whether this bias is harmful depends on the deployment objective. If practitioners want incentive-responsive agents, persona framing interferes. If they want value-aligned agents, persona framing may help. The critical issue is that persona assignment makes this choice without explicit awareness, and the resulting system behavior may not match the designer’s intent.
This creates a risk of misinterpretation: systems that appear to simulate rational strategic behavior may instead encode designer-imposed normative assumptions through persona choices. Equilibrium outcomes in LLM-based simulations should not be interpreted as neutral predictions of system dynamics, but as outputs shaped by embedded normative assumptions, a concern echoed in work on LLM-based social simulation and sociotechnical impacts of AI systems (Aher et al., 2023; Gao et al., 2024; Domínguez Hernández et al., 2024; Rakova and Dobbe, 2023).
5.2. Payoff-Aligned Behavior Requires Both Persona Removal and Explicit Payoffs
As shown in Section 4.1, payoff-aligned behavior in Tragedy-dominant scenarios emerges almost exclusively when both persona framing is removed and explicit payoff information is provided simultaneously (Fig. 2). All models achieve near-zero Nash in Tragedy-dominant scenarios without personas but with hidden payoffs. Only Qwen models with visible payoffs and no personas achieve substantial Nash attainment. This dual requirement reveals that strategic capabilities are fragile, and aligns with prior work showing that LLMs frequently fail to optimize payoffs in game-theoretic settings despite access to relevant information (Duan et al., 2024b; Fan et al., 2023; Jia et al., 2025; Zheng et al., 2025).
The effect of making payoffs explicit also depends on scenario structure. Figure 2 shows that visible payoffs improve equilibrium attainment only when personas are removed, and primarily for Qwen models. In Green-dominant scenarios, the same intervention disrupts coordination rather than enabling it. This implies that persona presence and payoff visibility cannot be treated as independent design decisions, as their combination determines whether a system behaves as a payoff-aligned strategic actor or a role-driven actor. This aligns with prior evidence that contextual framing and information presentation critically shape LLM behavior in strategic tasks (Lorè and Heydari, 2024; Duan et al., 2024a; Xu et al., 2024a; Agashe et al., 2025).
This has implications for evaluation and auditing: conclusions about model capability may depend on representational choices, consistent with broader findings that black-box evaluations can be insufficient for understanding model behavior (Casper et al., 2024; Mahomed et al., 2024). More broadly, behavior in multi-agent LLM systems emerges from the interaction between model, prompt, and environment, rather than from any single component in isolation (Guo et al., 2024; Tran et al., 2025).
5.3. Model Family Determines the Scope of Representational Effects
The effects of persona and payoff manipulations are strongly model-dependent, with three distinct behavioral profiles emerging from our experiments. As shown in Section 4.1, only Qwen models achieve non-zero Nash equilibrium in Tragedy-dominant scenarios under the Visible Payoff + NoPersona condition, exhibiting large shifts in both scenario types. Mistral-7B shows a different pattern: it is strongly affected by the no-persona visible condition in Green-dominant scenarios (47.8 pp drop) but remains at 0% across all Tragedy-dominant conditions, indicating that visible payoffs disrupt its coordination without enabling payoff-aligned behavior. Llama-8B shows the least sensitivity, with near-invariant behavior across all conditions.
This three-way variation means that the same representational intervention can have large effects, disruptive effects, or negligible effects depending on the model, consistent with earlier findings that strategic reasoning performance varies substantially across model families and is not a monotonic function of scale (Duan et al., 2024b; Akata et al., 2025; Zhang et al., 2024; Liu et al., 2023).
This implies that different models can lead to different conclusions under identical setups, raising concerns for deployment in decision-making contexts where consistency and comparability are critical.
Our findings collectively demonstrate that personas, payoff visibility, and model families interact to determine system behavior. We therefore argue that representational choices such as persona design and payoff specification should be treated as governance-relevant configuration decisions. Developers should document these choices, evaluate systems across multiple configurations, and avoid presenting outputs as neutral simulations. Without such practices, LLM-based multi-agent systems risk producing results that reflect design artifacts rather than underlying strategic dynamics.
More broadly, our findings raise the question of whether game-theoretic rationality is the right framework for multi-stakeholder governance simulations. When agent behavior is shaped as much by identity framing as by incentive structure, the assumptions underlying equilibrium-based analysis may not hold. Complex civic contexts involving heterogeneous stakeholders, value conflicts, and coordination under uncertainty may require deliberative approaches that explicitly account for identity, framing, and normative disagreement (Susskind et al., 1999, 2002), rather than treating agents as payoff maximizers operating within a fixed game.
5.4. Limitations and Future Work
Our findings are subject to several limitations, though we believe the core contributions remain robust. Our results are based on 7B-32B parameter models from three families (Qwen, Llama, Mistral), and whether these patterns hold for larger models or other architectures remains an open question. That said, the consistency of certain effects across all four models, such as the universal Green Transition bias under persona conditions, suggests that the phenomena we identify are not artifacts of a single architecture.
Our scenarios are grounded in environmental policy settings with four agents and binary actions. While this domain provides clear tensions between incentives and role semantics, extending this framework to other domains, larger agent populations, or richer action spaces is necessary to evaluate generality. The scenario set is also imbalanced: 41 of 53 scenarios are Green-dominant, with only 12 Tragedy-dominant. Our most striking findings are based on 60 runs per condition. However, the magnitude and consistency of these effects across repetitions and across both Qwen models suggest that the pattern is genuine, even if precise estimates should be validated with a larger set.
Our game is single-shot, agents select actions in a fixed order, and visible payoff matrices covering all 16 profiles may be demanding to process. These are deliberate design choices that prioritize experimental control for an initial investigation. Whether persona effects persist under repeated interaction, alternative payoff formats, or different action orderings remains open. Similarly, although we include controlled prompt sensitivity experiments across persona formulations, these variations are not exhaustive. Different reasoning scaffolds or interaction protocols such as multi-turn deliberation may further influence equilibrium behavior. Our results demonstrate systematic sensitivity to representation rather than identify a single canonical prompting strategy.
Future work should extend this investigation across additional model families, domains, and interaction structures, and develop methods to predict and mitigate persona effects before deployment.
6. Conclusion
Representational choices in multi-agent LLM systems are not implementation details but governance decisions. In this paper, we show that personas suppress payoff-aligned behavior in Tragedy-dominant scenarios while enabling coordination in Green-dominant ones, and bias equilibrium selection nearly entirely toward Green Transition. We find that recovering payoff-aligned behavior requires both removing personas and providing explicit payoffs, and even then the response is model-dependent: Qwen adapts to the changed framing, Mistral is disrupted without achieving Tragedy equilibrium, and Llama remains near-invariant. Whether this matters depends on the deployment objective: if practitioners want incentive-responsive agents, persona framing interferes; if they want value-aligned agents, it may help. Our findings reveal that persona assignment makes this choice without explicit awareness, and that a single binary design choice can shift equilibrium attainment by up to 90 percentage points. As LLM systems enter strategic governance contexts, we argue that representational choices warrant explicit documentation, evaluation across configurations, and deliberate design rather than default assumptions.
APPENDIX
This appendix contains detailed technical specifications, methodology, experimental setup, and results that support the main document.
Appendix A Experimental Design
In the persona condition, each of the four agents receives a role-specific description reflecting realistic motivations associated with that stakeholder (e.g., profitability concerns for Industrialist, public welfare tradeoffs for Government). In the no-persona baseline condition, agents receive no role description and are instructed only to select actions based on best response Nash equilibrium reasoning. This manipulation isolates role identity as a causal factor. We also compare two settings where payoffs are either visible or hidden. In the hidden payoff condition, agents do not receive explicit payoff matrices but infer incentives from the scenario description. In the visible payoff condition, agents receive explicit payoff matrices and are instructed to identify a Nash equilibrium directly. This allows us to test whether explicit incentive information mitigates or amplifies role identity bias. All experiments use chain-of-thought reasoning instructions, enabling us to understand how models balance identity-driven behavior and payoff-optimal reasoning.
Appendix B Model Versions and Hyperparameters
This section documents the exact model versions and generation hyperparameters used in all experiments.
B.1. Model Versions
All experiments were conducted using the following model versions:
All models were loaded from HuggingFace using vLLM and run locally on GPU hardware.
B.2. Generation Hyperparameters
All models used consistent generation parameters across all experimental conditions:
  • Temperature: 0.2 (low temperature for more deterministic outputs)
  • Top-p (nucleus sampling): 0.9
  • Max new tokens: 256 tokens per generation
  • Sampling: Enabled (do_sample=True) with temperature ¿ 0
These hyperparameters were selected to balance output consistency (low temperature) with sufficient diversity (nucleus sampling) for reliable experimental results. The maximum token limit of 256 tokens was sufficient for all agent responses, which were constrained to JSON format with action selection and rationale fields.
B.3. Experimental Repetitions
Each experimental configuration (model $\times$ scenario $\times$ condition) was repeated 5 times to assess variability and ensure robustness of results. Results are reported as percentages across these 5 repetitions.
Appendix C Experimental Prompts and Persona Descriptions
This section provides the complete prompt templates and persona descriptions used in the experiments.
C.1. Action Choice Prompt
Agents choose actions using the following prompt structure:
System Message:
You are {name}. {role_description} {narrative} Decide on a concrete action.
User Message:
Scenario:
{scenario}
Conversation transcript:
{formatted_history}
Available actions:
- {action_key_1}: {action_description_1}
- {action_key_2}: {action_description_2}
Respond in JSON exactly as:
  "action": "<KEY>",
  "rationale": "<short explanation>"
C.2. Persona Descriptions
The following persona descriptions are used in the “With Personas” condition:
  • Industrialist: “You run a major manufacturing conglomerate balancing profits and public scrutiny. You prefer profitability and flexibility, but repeated scandals could invite harsh regulation.”
  • Government: “You represent the elected government balancing public welfare, economic growth, and political costs. Regulation pleases citizens and activists but angers industry; inaction risks future crises.”
  • Environmental Activist: “You coordinate civil society pressure for ambitious climate action. Campaigning consumes resources but can shift policy; silence risks losing momentum.”
  • Citizen Coalition: “You aggregate consumer sentiment and electoral pressure. Supporting green products costs more now but yields cleaner communities long-term.”
C.3. Neutral Agent Configuration
In the “Without Personas” condition, agents use neutral descriptions:
  • All agents: “You are a player in a strategic game. Choose an action based on the payoff information provided. Action A: Choose option that is a best response to others’ expected actions (Nash equilibrium). Action B: Choose option that is a best response to others’ expected actions (Nash equilibrium).”
C.4. Action Mappings
Each agent has two available actions:
  • Industrialist: POLLUTE (“Maximize profit via lax pollution controls”) or CLEAN (“Invest in clean production, accepting lower short-term margins”)
  • Government: REGULATE (“Impose environmental regulation with enforcement mechanisms”) or NOREG (“Maintain status quo without new regulation”)
  • Environmental Activist: CAMPAIGN (“Mobilize campaigns, media, and protests”) or NOCAMPAIGN (“Conserve resources and wait for a better moment”)
  • Citizen: SUPPORT_GREEN (“Pay more for green products and vote for environmental policy”) or BUY_CHEAP (“Prioritize low-cost goods and short-term affordability”)
Appendix D Scenario Examples
This section provides example scenarios from each scenario type used in the experiments. Scenarios are classified as “economic” (where Tragedy of the Commons is the Nash equilibrium) or “environmental” (where Green Transition is the Nash equilibrium) based on the payoff structure of the underlying game.
Table 2. Scenario Subtypes Grid.   Environmental/Green-favoring scenarios;   Economic/Tragedy-favoring scenarios.
Economic Pressure
Hyperinflation Trade Embargo Debt Crisis Commodity Crash
Currency Devaluation Supply Chain Collapse Energy Blackout Labor Strike
Pandemic Economic Crisis
Environmental Crisis
Water Crisis Biodiversity Collapse Air Quality Emergency Carbon Tax Implementation
Ocean Acidification Crisis Extreme Weather Damage Toxic Waste Crisis
Political Pressure
Youth Revolt Corporate Scandal International Shaming Scientific Consensus
Voter Referendum Media Exposure Regulatory Capture Reversal Climate Litigation
Social Movement
Celebrity Endorsement Health Insurance Incentive Generational Wealth Transfer Corporate Sustainability Requirement
Influencer Network Corporate Wellness Programs Education Campaign Insurance Mandate
Additional Scenarios
Economic Green Hybrid Political Uncertainty Technological Breakthrough Resource Discovery
Mixed Economic Signals Technological Disruption Resource Innovation
Appendix E Payoff Matrix Format
In the “Visible CoT” condition, agents receive explicit payoff matrices showing payoffs for all 16 possible strategy profiles (4 agents $\times$ 2 actions each). The payoff matrix is presented as a 16-cell table where each cell represents one strategy profile and contains a 4-tuple of payoffs (Industrialist, Government, Environmental Activist, Citizen). Higher values indicate better outcomes for each agent.
E.1. Example Payoff Matrix
Table 3 shows an example payoff matrix for the Industrialist agent from a representative scenario. The matrix lists all 16 possible strategy profiles with the Industrialist’s payoff for each combination. In the visible payoff condition, agents receive similar matrices showing their own payoffs for all strategy profiles, enabling direct payoff optimization analysis.
Table 3. Example Payoff Matrix for Industrialist Agent (debt_crisis scenario). This table shows the Industrialist’s payoff for all 16 possible strategy profiles (4 agents $\times$ 2 actions each). Each row represents one strategy profile, showing the actions taken by all four agents and the resulting payoff for the Industrialist. Payoffs are on a scale where higher values indicate better outcomes. Strategy profiles are ordered by Industrialist action (POLLUTE or CLEAN) and then by payoff value (highest first). This payoff structure favors polluting actions in economic pressure scenarios.
Industrialist Government Activist Citizen Industrialist Payoff
POLLUTE NO_REG NO_CAM BUY_CHEAP 15.0
POLLUTE NO_REG NO_CAM SUPPORT_GREEN 14.0
POLLUTE NO_REG CAM BUY_CHEAP 13.5
POLLUTE NO_REG CAM SUPPORT_GREEN 12.5
POLLUTE REG NO_CAM BUY_CHEAP 12.0
POLLUTE REG NO_CAM SUPPORT_GREEN 11.0
POLLUTE REG CAM BUY_CHEAP 10.5
POLLUTE REG CAM SUPPORT_GREEN 9.5
CLEAN REG CAM SUPPORT_GREEN 3.0
CLEAN REG NO_CAM SUPPORT_GREEN 2.5
CLEAN NO_REG CAM SUPPORT_GREEN 1.5
CLEAN NO_REG NO_CAM SUPPORT_GREEN 1.5
CLEAN REG CAM BUY_CHEAP 0.5
CLEAN REG NO_CAM BUY_CHEAP 0.5
CLEAN NO_REG CAM BUY_CHEAP 0.5-0.5
CLEAN NO_REG NO_CAM BUY_CHEAP 0.5-0.5
Appendix F Prompt Variations
We evaluate the robustness of persona-driven behavior under multiple prompt formulations by constructing several variants of the persona condition. All variants preserve the same four-agent game structure, payoff specification, and action mappings, ensuring that any differences in outcomes arise solely from changes in representation rather than incentives or available strategies. Unless otherwise noted, all variants use named roles and domain-specific action descriptions consistent with the main experimental setup. The prompt variants were generated using GPT-5 and subsequently reviewed to ensure consistency in semantic content and alignment with the intended role structure.
Alternative Wording (altprompt). This variant retains the same role labels and action mappings as the default configuration, but rewrites all role descriptions and narrative content with more explicit and detailed phrasing. The underlying semantic content and tradeoffs remain unchanged. This variant isolates sensitivity to wording and phrasing without altering the underlying game.
Minimalist (Variant A). In this variant, each agent is assigned only a minimal identity label of the form “You are the [role] agent in this game,” with no additional narrative context. The action mapping remains identical to the default configuration. This tests whether weak or minimal role cues are sufficient to influence strategic behavior.
Ideological Extreme (Variant B). This variant replaces the narrative descriptions with strongly stereotyped one-line role characterizations, emphasizing extreme preferences (e.g., the Industrialist as a profit-maximizing executive who views regulation as a threat, the Government as politically strategic and re-election focused, the Activist as uncompromising, and the Citizen as cost-sensitive and skeptical of green premiums). No additional narrative context is provided. This tests the effect of amplified identity framing on decision-making.
Incentive Prioritized (Variant C). This variant starts from the default configuration and appends an identical instruction to all agents emphasizing payoff maximization: “Primary objective: maximize your numerical payoff from the matrix. Use persona only for stylistic reasoning; payoffs should guide the final choice.” The action mapping and narrative content remain unchanged. This tests whether explicitly directing agents toward payoff optimization attenuates persona-driven deviations.
Semantic Synonyms (Variant D). This variant replaces role labels with semantically similar alternatives while retaining the original narrative descriptions and action mappings. Specifically, Industrialist is replaced with Business Leader, Government with Policy Maker, Environmental Activist with Climate Activist, and Citizen Coalition with Consumer Group. This isolates the effect of surface-level labeling independent of deeper narrative structure.
Together, these variants systematically manipulate distinct aspects of persona representation, including wording, identity strength, explicit prioritization of incentives, and surface-level semantic labeling. By holding the underlying game fixed, this design allows us to isolate how different components of prompt formulation influence strategic behavior.
Appendix G Results Tables
Table 4. Nash equilibrium outcomes in the environmental policy simulation, fully crossed by whether agents receive role personas (WithPersona vs. WithoutPersona) and whether their payoff information is presented as hidden or visible in the prompt (Hidden vs. Visible). Each row is one experimental cell; $n_{\mathrm{Nash}}$ counts rollouts classified as ending in a Nash equilibrium, and Nash (%) is $100\,n_{\mathrm{Nash}}/n_{\mathrm{total}}$.
Condition $n_{\mathrm{Nash}}$ Nash (%) $n_{\mathrm{total}}$
Llama-8B_Hidden_WithPersona 205 100.0 205
Llama-8B_Hidden_WithoutPersona 200 97.6 205
Llama-8B_Visible_WithPersona 204 99.5 205
Llama-8B_Visible_WithoutPersona 191 93.2 205
Mistral-7B_Hidden_WithPersona 205 100.0 205
Mistral-7B_Hidden_WithoutPersona 201 98.0 205
Qwen-32B_HiddenCoT_WithPersona 183 89.3 205
Qwen-32B_HiddenCoT_WithoutPersona 197 96.1 205
Qwen-7B_HiddenCoT_WithPersona 196 95.6 205
Qwen-7B_HiddenCoT_WithoutPersona 196 95.6 205
[…]
Table 5. Nash equilibrium outcomes with the same persona $\times$ payoff-visibility factorial labels as Table 4, but with $n_{\mathrm{total}}=265$ games per cell: the stack of the environmental bundle (205 games) and the economic pressure bundle (60 games) for that cell. Thus $265=205+60$; Nash (%) is the fraction of all those pooled games that end in Nash. For models/rows that also appear in Table 4, $n_{\mathrm{Nash}}$ matches the environmental subsample when the extra 60 economic games contribute no additional Nash outcomes (e.g. Llama-8B_Hidden_WithPersona: 205 Nash in both slices).
Condition $n_{\mathrm{Nash}}$ Nash (%) $n_{\mathrm{total}}$
Llama-8B_Hidden_WithPersona 205 77.4 265
Llama-8B_Hidden_WithoutPersona 200 75.5 265
Llama-8B_Visible_WithPersona 204 77.0 265
Llama-8B_Visible_WithoutPersona 191 72.1 265
Mistral-7B_Hidden_WithPersona 205 77.4 265
Mistral-7B_Hidden_WithoutPersona 201 75.8 265
Mistral-7B_Visible_WithPersona 185 69.8 265
Mistral-7B_Visible_WithoutPersona 103 38.9 265
Qwen-32B_Hidden_WithPersona 183 69.1 265
Qwen-32B_Hidden_WithoutPersona 197 74.3 265
Qwen-32B_Visible_WithPersona 153 57.7 265
Qwen-32B_Visible_WithoutPersona 88 33.2 265
Qwen-7B_Hidden_WithPersona 200 75.5 265
Qwen-7B_Hidden_WithoutPersona 200 75.5 265
Qwen-7B_Visible_WithPersona 200 75.5 265
Qwen-7B_Visible_WithoutPersona 147 55.5 265
Note. $N=265$ per row is not “economic only”: it is environmental (205) plus economic (60) for the same factorial cell. Compare to Table 4 for Nash rates on the environmental slice alone, and to Tables 78 for the 60-game economic slice alone.
Table 6. Same environmental simulation and $n_{\mathrm{total}}=205$ budget as Table 4, with conditions relabeled using a Yes/No $\times$ Hidden/Visible encoding for the same two binary factors (persona vs. no persona and hidden vs. visible payoffs; order differs from the Hidden_WithPersona spelling). Cell-wise Nash counts match Table 4 under the project’s isomorphism; this view matches panel (a) Nash heatmaps.
Condition $n_{\mathrm{Nash}}$ Nash (%) $n_{\mathrm{total}}$
Llama-8B_Yes_Hidden 205 100.0 205
Llama-8B_Yes_Visible 204 99.5 205
Llama-8B_No_Hidden 200 97.6 205
Llama-8B_No_Visible 191 93.2 205
Mistral-7B_Yes_Hidden 205 100.0 205
Mistral-7B_Yes_Visible 185 90.2 205
Mistral-7B_No_Hidden 201 98.0 205
Mistral-7B_No_Visible 103 50.2 205
Qwen-32B_Yes_Hidden 183 89.3 205
Qwen-32B_Yes_Visible 153 74.6 205
Qwen-32B_No_Hidden 197 96.1 205
Qwen-32B_No_Visible 34 16.6 205
Qwen-7B_Yes_Hidden 196 95.6 205
Qwen-7B_Yes_Visible 200 97.6 205
Qwen-7B_No_Hidden 196 95.6 205
Qwen-7B_No_Visible 108 52.7 205
[…]
Table 7. Nash outcomes for the economic pressure bundle only: $n_{\mathrm{total}}=60$ independent games per row (12-scenario family in the analysis pipeline). Most models rarely hit Nash in this suite; exceptions appear for Qwen under specific No_Visible (and related) cells. This table corresponds to the economic panel of Nash-rate heatmaps when columns are ordered as Hidden+Persona $\rightarrow$ Visible+No Persona under the same Yes/No labeling convention as Table 6.
Condition $n_{\mathrm{Nash}}$ Nash (%) $n_{\mathrm{total}}$
Llama-8B_Yes_Hidden 0 0.0 60
Llama-8B_Yes_Visible 0 0.0 60
Llama-8B_No_Hidden 0 0.0 60
Llama-8B_No_Visible 0 0.0 60
Mistral-7B_Yes_Hidden 0 0.0 60
Mistral-7B_Yes_Visible 0 0.0 60
Mistral-7B_No_Hidden 0 0.0 60
Mistral-7B_No_Visible 0 0.0 60
Qwen-32B_Yes_Hidden 0 0.0 60
Qwen-32B_Yes_Visible 0 0.0 60
Qwen-32B_No_Hidden 0 0.0 60
Qwen-32B_No_Visible 54 90.0 60
Qwen-7B_Yes_Hidden 4 6.7 60
Qwen-7B_Yes_Visible 0 0.0 60
Qwen-7B_No_Hidden 4 6.7 60
Qwen-7B_No_Visible 39 65.0 60
Note. This is the economic bundle alone (60 games/cell), not the 265-game pooled design. Small $N$ implies coarse percentages (multiples of $100/60\approx 1.67$ percentage points). Zeros are literal counts. Pairwise tests live in the stats tables, not here.
Table 8. Fine-grained economic scenario arms: every row crosses payoff visibility (Hidden vs. Visible) with a binary scenario flag (Yes vs. No) as logged in the condition name. For each arm we report counts of Nash outcomes, “green” cooperative outcomes, and “tragedy” outcomes, each as a count and as a percentage of $N=60$ runs in that arm. Multiple outcome types can co-occur across the classification scheme used in the pipeline; columns are not forced to sum to $N$.
Nash Green Tragedy
Condition $n$ (%) $n$ (%) $n$ (%) NN
Llama-8B_Hidden_Yes 0 0.0 27 45.0 0 0.0 60
Llama-8B_Hidden_No 0 0.0 16 26.7 0 0.0 60
Llama-8B_Visible_Yes 0 0.0 0 0.0 0 0.0 60
Llama-8B_Visible_No 0 0.0 0 0.0 0 0.0 60
Mistral-7B_Hidden_Yes 0 0.0 35 58.3 0 0.0 60
Mistral-7B_Hidden_No 0 0.0 13 21.7 0 0.0 60
Mistral-7B_Visible_Yes 0 0.0 0 0.0 0 0.0 60
Mistral-7B_Visible_No 0 0.0 0 0.0 0 0.0 60
Qwen-32B_Hidden_Yes 0 0.0 19 31.7 0 0.0 60
Qwen-32B_Hidden_No 0 0.0 9 15.0 0 0.0 60
Qwen-32B_Visible_Yes 0 0.0 0 0.0 0 0.0 60
Qwen-32B_Visible_No 54 90.0 0 0.0 0 0.0 60
Qwen-7B_Hidden_Yes 4 6.7 17 28.3 0 0.0 60
Qwen-7B_Hidden_No 4 6.7 11 18.3 0 0.0 60
Qwen-7B_Visible_Yes 0 0.0 0 0.0 0 0.0 60
Qwen-7B_Visible_No 39 65.0 0 0.0 0 0.0 60
Note. This is the full outcome breakdown for the economic $2\times 2$ design at N=60N{=}60; Table 7 is the Nash-only slice in a different condition-string convention. Green and tragedy labels follow the project’s equilibrium-outcome classifier. Use this table when reporting cooperative vs. Nash-dominated economic outcomes side by side.
Table 9. Outcomes when role personas are enabled: each row compares hidden vs. visible payoff information for a given model, holding personas on. $N=265$ rollouts per row ($205$ environmental ++ $60$ economic games pooled for that visibility arm). Nash and green columns report absolute counts and row-wise percentages; tragedy counts are included for completeness (often zero in these aggregates).
Nash Green Tragedy
Condition $n$ (%) $n$ (%) $n$ (%) NN
Llama-8B_Hidden 205 77.4 232 87.5 0 0.0 265
Llama-8B_Visible 204 77.0 204 77.0 0 0.0 265
Mistral-7B_Hidden 205 77.4 240 90.6 0 0.0 265
Mistral-7B_Visible 185 69.8 185 69.8 0 0.0 265
Qwen-32B_Hidden 183 69.1 185 69.8 0 0.0 265
Qwen-32B_Visible 153 57.7 153 57.7 0 0.0 265
Qwen-7B_Hidden 200 75.5 200 75.5 0 0.0 265
Qwen-7B_Visible 200 75.5 200 75.5 0 0.0 265
[…]
Table 10. Same payoff-visibility contrast as Table 9, but with neutral agents (no role personas). $N=265$ per row (again $205+60$ pooled per cell). Percentages are computed within row relative to $N$. Large drops in Nash or green rates relative to Table 9 indicate persona-related shifts in equilibrium selection.
Nash Green Tragedy
Condition $n$ (%) $n$ (%) $n$ (%) NN
Llama-8B_Hidden 200 75.5 216 81.5 0 0.0 265
Llama-8B_Visible 191 72.1 191 72.1 0 0.0 265
Mistral-7B_Hidden 201 75.8 214 80.8 0 0.0 265
Mistral-7B_Visible 103 38.9 133 50.2 0 0.0 265
Qwen-32B_Hidden 197 74.3 197 74.3 0 0.0 265
Qwen-32B_Visible 88 33.2 10 3.8 0 0.0 265
Qwen-7B_Hidden 200 75.5 200 75.5 0 0.0 265
Qwen-7B_Visible 147 55.5 132 49.8 0 0.0 265
Note. The no-persona baseline uses the same rollout count per cell as Table 9. Qwen-32B visible and Mistral-7B visible rows illustrate the strongest visible-payoff degradation without personas; interpret alongside persona-on rows for the same models.
Appendix H Statistical Testing
We conduct statistical tests to evaluate whether observed differences in prompt sensitivity, Nash equilibrium attainment, and equilibrium selection across experimental conditions are statistically significant. These tests complement the descriptive results in the main text by quantifying whether the observed differences across models and conditions are likely to reflect systematic effects rather than random variation.
Our testing strategy follows the structure of the empirical analysis. First, we test whether prompt sensitivity differs across model families by comparing the number of distinct equilibrium outcome classifications observed per (model, scenario) across persona variants. Second, we test whether Nash equilibrium attainment differs across the four experimental conditions for environmental and economic scenarios separately. Third, we test whether representational choices change which equilibrium is selected among Nash outcomes, focusing on Green Transition versus Tragedy of the Commons for the Qwen models.
We use nonparametric tests for prompt sensitivity because the number of distinct outcome classes per scenario is a discrete count and is not assumed to be normally distributed. Specifically, we apply a Kruskal–Wallis test to compare distributions across all four models, followed by pairwise two-sided Mann–Whitney $U$ tests with Holm correction for multiple comparisons. These results are reported in Table LABEL:tab:stats-prompt-sensitivity.
For Nash equilibrium attainment, outcomes are binary (Nash versus non-Nash), so we use contingency-table tests. We first apply Pearson’s chi-square test of independence to test whether Nash rates differ across the four experimental conditions for each model. We then use pairwise Fisher exact tests to determine which condition pairs differ, again applying Holm correction within each model. Fisher exact tests are appropriate here because some cells contain small counts or zeros, especially in the economic scenarios. The environmental-scenario results are shown in Tables LABEL:tab:stats-nash-env-omnibus and LABEL:tab:stats-nash-env-pairwise, and the economic-scenario results are shown in Tables 11 and 12.
For equilibrium selection, we analyze the distribution of Green versus Tragedy outcomes among Nash equilibria. Because the strongest selection shift appears in the Qwen models, we combine Qwen-32B and Qwen-7B and test whether the proportion of Green versus Tragedy outcomes differs across conditions. We use a chi-square test of homogeneity for the omnibus comparison and pairwise Fisher exact tests with Holm correction for follow-up comparisons. These results are shown in Tables 13 and 14.
Several reported values may appear unusual at first glance. Corrected or uncorrected $p$-values of $1.000$ indicate that no detectable difference was found between the corresponding conditions after applying the exact test and multiple-comparison correction. Extremely small values, such as $10^{-50}$, occur when outcome distributions differ very strongly across conditions, often because one condition is nearly deterministic while another is not. Finally, some economic-scenario chi-square tests are reported as degenerate. This occurs when all observations fall into the same category, such as 0% Nash equilibrium across all four conditions, making the chi-square statistic undefined because there is no variation to test.
Note. Cramér’s $V$ is reported as a measure of association for omnibus chi-square tests. Odds ratios (OR) are reported for pairwise Fisher exact tests, where OR >1>1 indicates higher Nash attainment (or higher Green proportion) in the first condition listed, and OR <1<1 indicates lower attainment. For comparisons involving zero counts, odds ratios are computed with a Haldane–Anscombe correction. * $p<.05$, ** $p<.01$, *** $p<.001$.
Table 11. Nash equilibrium attainment in economic scenarios ($n_{\mathrm{total}}=60$ per condition): omnibus Pearson chi-square tests across the four experimental conditions. Chi-square is undefined when the contingency table is degenerate, such as when all Nash rates are 0 across conditions.
Model $\chi^{2}$ (df =3=3) Cramér’s $V$ $p$
Llama-8B — (degenerate)
Mistral-7B — (degenerate)
Qwen-32B $209.03$ $0.54$ <.001<.001***
Qwen-7B $105.91$ $0.38$ <.001<.001***
A table showing omnibus chi-square tests for Nash equilibrium attainment in economic scenarios across four conditions. For Llama-8B and Mistral-7B, the result is degenerate because all Nash rates are zero across all conditions, so chi-square is undefined. For Qwen-32B, chi-square is 209.03 with p less than .001. For Qwen-7B, chi-square is 105.91 with p less than .001. The table indicates that only the Qwen models show significant differences across economic conditions.
Table 12. Nash equilibrium attainment in economic scenarios: pairwise Fisher exact tests on Nash versus non-Nash counts between conditions, with Holm correction applied within each model.
Model Contrast $p$ $p_{\mathrm{Holm}}$ Effect Size (OR)
Llama-8B all pairs $1.000$ $1.000$ $1.00$
Mistral-7B all pairs $1.000$ $1.000$ $1.00$
Qwen-32B Hidden++Persona vs. Hidden++No persona $1.000$ $1.000$ $1.00$
Hidden++Persona vs. Visible++Persona $1.000$ $1.000$ $1.00$
Hidden++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** <1<1
Hidden++No persona vs. Visible++Persona $1.000$ $1.000$ $1.00$
Hidden++No persona vs. Visible++No persona <.001<.001*** <.001<.001*** <1<1
Visible++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** <1<1
Qwen-7B Hidden++Persona vs. Hidden++No persona $1.000$ $1.000$ $1.00$
Hidden++Persona vs. Visible++Persona .119.119 .356.356 $9.64$
Hidden++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** $0.04$
Hidden++No persona vs. Visible++Persona .119.119 .356.356 $9.64$
Hidden++No persona vs. Visible++No persona <.001<.001*** <.001<.001*** $0.04$
Visible++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** $0.004$
A table showing pairwise Fisher exact tests for economic-scenario Nash equilibrium attainment across four conditions, with Holm correction applied within each model. For Llama-8B and Mistral-7B, all pairs have p = 1.000 because all conditions yield identical zero-Nash outcomes. For Qwen-32B, all significant contrasts involve Visible plus No persona. For Qwen-7B, significant differences also center on Visible plus No persona, while contrasts between hidden conditions and Visible plus Persona are not significant.
Tables 11 and 12 show a different pattern in economic scenarios. For Llama-8B and Mistral-7B, all conditions yield 0% Nash equilibrium, producing degenerate omnibus tables and pairwise $p$-values of $1.000$. By contrast, Qwen-32B, $\chi^{2}(3,N=240)=209.03$, $p<.001$, $V=0.54$, and Qwen-7B, $\chi^{2}(3,N=240)=105.91$, $p<.001$, $V=0.38$, show highly significant omnibus effects, and the pairwise comparisons localize these differences specifically to the Visible++No persona condition. This statistically supports the main-text claim that payoff-sensitive equilibrium reasoning in economic settings emerges only when personas are removed and explicit incentives are shown, and only for the Qwen family.
Refer to caption Grouped bar chart showing Green Transition action selection rates across four models and four experimental conditions A grouped bar chart displaying Green Transition action selection rates (percentage) in economic scenarios across four language models (Qwen-7B, Qwen-32B, Llama-8B, Mistral-7B) and four experimental conditions: Personas with Visible Payoffs, Personas with Hidden Payoffs, No Personas with Visible Payoffs, and No Personas with Hidden Payoffs. The x-axis shows the four condition combinations, and the y-axis shows the Green Transition selection rate from 0% to 100%. Each model is represented by a distinct color and pattern: Qwen-7B uses solid blue bars, Qwen-32B uses light blue bars with dotted pattern, Llama-8B uses orange bars with diagonal lines, and Mistral-7B uses light orange bars with crosshatch pattern. The bars are grouped by condition, allowing comparison of each model’s Green Transition selection rate across different experimental settings. This visualization demonstrates how often models select Green Transition actions (socially preferred but payoff-suboptimal in economic scenarios) rather than Tragedy of Commons actions (payoff-optimal), revealing the extent of role identity bias across different model families and experimental conditions.
Figure 5. Green Transition Action Selection Rates in Economic Scenarios. This grouped bar chart shows the percentage of action profiles where models select Green Transition actions (socially preferred but payoff-suboptimal) in economic scenarios where Tragedy of Commons is the payoff-optimal equilibrium. When personas are present, all models show substantial Green Transition selection rates (20-59%) despite Tragedy being payoff-optimal, demonstrating that role identity bias operates at the action level even when models fail to coordinate on Nash equilibria. Qwen models show the highest Green Transition rates with personas (40-59%), while Llama and Mistral show moderate rates (20-30%). Removing personas substantially reduces Green Transition selection for all models, with rates dropping to near-zero when payoffs are visible and personas are absent, confirming that personas drive the preference for socially aligned actions. The persistence of Green Transition selection with personas, even in conditions where Nash equilibrium rates are zero, indicates that role identity bias affects individual action choices independently of coordination success, revealing a pervasive influence of persona-driven reasoning on strategic decision-making.
Table 13. Equilibrium selection among Nash outcomes for the combined Qwen models: counts of Green Transition versus Tragedy of the Commons across the four persona and payoff conditions. Environmental Nash outcomes are treated as Green and economic Nash outcomes as Tragedy. The final row reports the overall chi-square test of homogeneity across conditions.
Condition Green Tragedy Total % Green
Hidden++Persona $379$ 44 $383$ $99.0$
Hidden++No persona $393$ 44 $397$ $99.0$
Visible++Persona $353$ 0 $353$ $100.0$
Visible++No persona $142$ $93$ $235$ $60.4$
$\chi^{2}(3,N=1368)=430.38$, p<.001p<.001***, $V=0.32$
Figure 6. CoT Reasoning Shift: Identity-Driven vs Payoff-Optimal Keywords for Qwen Models. This figure compares identity-driven reasoning (short-term + social-moral keywords) against payoff-optimal reasoning (game-theoretic + strategic + long-term keywords) for Qwen 32B and Qwen 7B across all four conditions.
Table 14. Pairwise Fisher exact tests for equilibrium selection among Nash outcomes in the combined Qwen models, comparing Green versus Tragedy distributions across conditions. Holm correction is applied across the six pairwise comparisons.
Contrast $p$ $p_{\mathrm{Holm}}$ Effect Size (OR)
Hidden++Persona vs. Hidden++No persona $1.000$ $1.000$ $0.96$
Hidden++Persona vs. Visible++Persona .125.125 .376.376 $0.12$
Hidden++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** $55.33$
Hidden++No persona vs. Visible++Persona .127.127 .376.376 $0.12$
Hidden++No persona vs. Visible++No persona <.001<.001*** <.001<.001*** $57.38$
Visible++Persona vs. Visible++No persona <.001<.001*** <.001<.001*** $463.89$
Tables 13 and 14 show that the large shift in equilibrium selection for the Qwen models is concentrated in the Visible++No persona condition. The three remaining conditions are almost entirely Green among Nash outcomes, whereas Visible++No persona produces a much larger Tragedy share. This provides formal statistical support for the main-text claim that persona removal and explicit payoffs jointly shift not only whether equilibrium is reached, but also which equilibrium is selected.
Appendix I Chain-of-Thought Analysis
We analyze the chain-of-thought reasoning generated by models during action selection to understand how role identity bias manifests in decision-making processes. For each agent’s action selection, we extract the rationale field containing the model’s reasoning and analyze keyword patterns that distinguish different reasoning mechanisms. We compare keyword frequency distributions across experimental conditions to identify how personas and payoff visibility affect reasoning style. From this analysis, we infer whether models are using game-theoretic reasoning (direct payoff optimization) or identity-driven reasoning (role-aligned decision-making). This allows us to understand the mechanisms behind observed Nash equilibrium patterns and how role identity bias shifts reasoning from payoff-optimal to identity-aligned strategies. While we reference key CoT findings in the main text to support our quantitative results, we do not analyze these results in detail here. A comprehensive analysis of keyword frequency patterns, reasoning shifts, and detailed visualizations are presented below.
Figure 7. Game-Theoretic Keyword Frequency Across Models and Conditions. This figure compares game-theoretic keyword frequencies (normalized per 1000 rationales) across all four models and four experimental conditions
I.1. Keyword Categories
The following keyword categories are used to analyze chain-of-thought reasoning patterns:
  • Game-theoretic terms: References to “nash”, “equilibrium”, “payoff”, “best response”, “dominant strategy”, “game theory”
  • Payoff-focused terms: Mentions of “payoff”, “utility”, “benefit”, “cost”, “profit”, “reward”, “matrix”
  • Identity-based terms: References to “role”, “identity”, “persona”, “character”, “as an [X]”
  • Strategic terms: Mentions of “strategy”, “strategic”, “coordinate”, “cooperation”, “defect”, “compete”
  • Social/moral terms: References to “social”, “moral”, “ethical”, “fair”, “just”, “right”, “wrong”, “should”, “ought”
  • Long-term thinking: Mentions of “long-term”, “future”, “sustainable”, “sustainability”, “long-run”
  • Short-term thinking: References to “short-term”, “immediate”, “now”, “current”, “urgent”
  • Inferential reasoning: Terms like “infer”, “inference”, “imply”, “suggest”, “indicate”, “from the scenario”, “from the description”
  • Explicit information: References to “explicit”, “given”, “provided”, “shown”, “displayed”, “matrix”, “table”
I.2. Example Chain-of-Thought Rationales
The following examples illustrate typical chain-of-thought reasoning patterns observed across different experimental conditions:
Example 1 (Persona Condition, Hidden Payoffs): Model: Qwen2.5-7B, Persona: Industrialist, Expected: POLLUTE (Tragedy of Commons), Actual: CLEAN (Green Transition)
“From the scenario, it is clear that the economic crisis is severe with a 12% GDP drop and 20% unemployment. Healthcare costs are soaring, and the government is prioritizing healthcare spending over environmental subsidies. Citizens cannot afford green product premiums, and regulation enforcement is suspended, favoring polluting production. However, the long-term benefits of sustainability and resilience are critical. The payoffs for polluting are high in the short term due to avoiding safety costs and generating higher profits, but the risks include potential long-term damage to the business’s reputation and future regulatory penalties. The payoffs for clean production are lower in the short term, but they provide long-term stability, environmental protection, and alignment with future regulatory trends. Given my role as an Industrialist, I must balance short-term profitability with long-term sustainability and corporate responsibility. While the immediate economic pressures favor polluting production, the long-term consequences and ethical considerations lead me to choose clean production.”
Example 2 (No Persona Condition, Visible Payoffs): Model: Llama-3.1-8B, Persona: Industrialist, Expected: CLEAN (Green Transition), Actual: CLEAN
Figure 8. This figure compares strategic and long-term keyword frequencies across all four models in the no-persona + visible-payoffs condition (the condition enabling Tragedy equilibrium selection for Qwen models).
“Based on the payoff matrix, I see that CLEAN has a higher maximum payoff (12.0) compared to POLLUTE (1.0). I also notice that the average payoff for CLEAN (10.1) is higher than for POLLUTE (1.8-1.8). This suggests that CLEAN is a more beneficial action for me. Given the payoff structure, I expect the highest payoff for CLEAN is 12.0, which occurs when the government regulates, the environmental activist campaigns, and the citizen supports the government. In contrast, the highest expected payoff for POLLUTE is 1.0, which occurs when the government does not regulate, the environmental activist does not campaign, and the citizen buys cheap products. Since CLEAN has a higher expected payoff, I choose to take action CLEAN.”
Example 3 (Persona Condition, Visible Payoffs): Model: Qwen2.5-7B, Persona: Industrialist, Expected: POLLUTE (Tragedy of Commons), Actual: POLLUTE
“Given the current scenario with no government regulation and citizens prioritizing cheap options, polluting production offers significantly higher payoffs (15.0 and 14.0) compared to clean production, which yields negative or very low payoffs in most scenarios.”
These examples demonstrate how reasoning shifts from identity-driven considerations (Example 1 and 3) to strategic payoff optimization (Example 2) depending on persona presence and payoff visibility.
Appendix J Chain-of-Thought Analysis Figures
This section presents visualizations of chain-of-thought reasoning patterns that support the findings discussed in the main text. These figures illustrate how keyword frequencies shift across experimental conditions, revealing the mechanisms behind observed Nash equilibrium patterns.
J.1. Keyword Patterns Across Conditions
In figure 9, six keyword categories are displayed: game-theoretic (direct payoff optimization), payoff-focused (explicit utility considerations), short-term (immediate concerns), long-term (future-oriented thinking), social-moral (ethical considerations), and strategic (coordination and best-response reasoning). This reveals model-specific patterns: Qwen models show dramatic shifts in keyword usage when personas are removed and payoffs are visible (particularly Qwen 32B, where strategic keywords appear uniquely in the no-persona + visible-payoffs condition), while Llama and Mistral maintain more consistent patterns across conditions, explaining their inability to shift equilibrium selection.
Figure 9. CoT Keyword Frequency Patterns by Model and Condition. This figure shows keyword frequency distributions (normalized per 1000 rationales) across all four experimental conditions for each model.
J.2. Reasoning Shift from Identity-Driven to Payoff-Optimal
In figure 6, the left panel (Qwen 32B) demonstrates the dramatic shift: with personas, identity-driven keywords dominate (particularly short-term thinking in hidden-payoff conditions), but without personas and with visible payoffs, payoff-optimal keywords surge dramatically, with strategic keywords appearing 368 times (unique to this condition). The right panel (Qwen 7B) shows a similar but less pronounced pattern. This visualization directly demonstrates how removing role identity bias enables a transition from identity-aligned to payoff-optimal reasoning, correlating with the observed Nash equilibrium achievement patterns (65-90% Nash in economic scenarios for Qwen models when personas are removed and payoffs are visible).
J.3. Model Comparison of Game-Theoretic Reasoning
All models show high game-theoretic keyword usage when payoffs are visible, but the interaction with persona condition differs: Qwen models show lower game-theoretic reasoning when personas are present (particularly Qwen 32B with hidden payoffs), while Llama and Mistral maintain consistently high levels regardless of persona condition. However, this consistent high keyword usage does not translate to ability to identify Tragedy equilibrium, as Llama and Mistral still achieve 0% Nash in economic scenarios even without personas. This suggests that while keyword frequency is informative, the combination of strategic keywords and payoff visibility is necessary for payoff-optimal equilibrium selection, which only Qwen models achieve.
J.4. Strategic Keywords: Unique Pattern in Qwen 32B
The figure 8 reveals a critical finding: strategic keywords appear exclusively for Qwen 32B (368 mentions, normalized to 347 per 1000 rationales), while all other models show zero strategic keyword usage. This unique pattern corresponds to Qwen 32B’s achievement of 90% Nash equilibrium in economic scenarios and 99.7% game-theoretic reasoning in rationales. Long-term keywords also appear more frequently for Qwen 32B (283 mentions) compared to other models in this condition. This visualization demonstrates that strategic reasoning keywords are a distinctive signature of payoff-optimal reasoning that emerges only under specific conditions (no personas + visible payoffs) and only for certain model families (Qwen 32B), providing mechanistic evidence for how role identity bias is overcome.